1

”

A Complete Bibliography of Publications in Journal of Open Research Software

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.04

Title word cross-reference

1 [DB17]. 2 [DB15]. 3 [ML16, WAH^{+}16]. ^{14} [LO16]. D [HH17]. N [HH17, Han22, vGvdB16]. t [Kel20].

-D [WAH^{+}16]. -dimensional [Han22]. -gram [vGvdB16]. -Tests [Kel20].

3 [LD19, MOM21]. 3.0 [NMA^{+}18]. 3D [BR21].

ABC [NS20]. ABC-OCT [NS20]. Abstractions [DBB^{+}14]. Accelerate [VHT^{+}19]. Accelerators [AC17]. Access [PdV21]. Accessible [VBAF16]. Accuracy [TS21]. Accurate [TS14]. Acoustic [BBJ17]. Acquisition [TDX^{+}20, Wag17].

Action [RD22]. Adaguc [PdV21]. Adaguc-Server [PdV21]. Adapting [CB16b]. ADAT [CFC^{+}17]. Additive [Sto16]. Address [Kat14]. AddressingHistory [OHM14]. Adjacency [XXX16]. Adopting [CB16b]. Adults [ME14]. Advanced [HFM^{+}21, MHE^{+}18, PM17]. Affect [Gir14, PDR^{+}17]. Affinity [XXX16]. Affordable [TS14]. Age [LO16]. Agent [KFV20]. Agent-Based [KFV20]. Alcohol [LPG^{+}18]. Algebraic [WS15]. Algorithm [LVSF16, Raß20]. Algorithmic [Grö20]. Algorithms [LW20, TS21]. Alida [PM17]. American [ES21, Kim17]. AMR [DV14]. Analysing [SM14a, vGFvNvA19]. Analysis [AL16, BVL18, BHS18, CRL21, CZZ19, CKDG20, Daw16a, ES21, FLMR16a, Far16, FCY20, FSBK18, FS16, GDP18, HH15, Han21, Han22, Law17, MGMP16, OK20, PM17, PDR^{+}17, RD22, Sta21, TLR21, Wag17, WL13, YL17, ZO16, vGFvNvA19, EBFS17, FSV^{+}13, PFLG21]. Analytical [ZO16]. Analytics [Gov16]. Analyze [BS22, KWV^{+}20]. Analyzing [FLMR16b]. Android [LPG^{+}18, Luk21]. Animal [BBJ17]. Annotation [Gir14, FPF14]. Annotator [HYG16]. Anterior [RD22]. API [CSR^{+}23, MBA19, Moo21, SGPHD^{+}17]. App [Luk21]. Application [BS22, EML21, Far16, GPP^{+}14, Pas16, RML17, SPG^{+}15]. Applications [DVS20, HFM^{+}21, MA13, ML16, PM17]. Apply [Haz14]. Approach [RLWP16]. Approximation [TCD^{+}22]. ARBTools [WKC19]. Architectures [DBB^{+}14, HFM^{+}21]. area [GRZ13]. ARM [HC16]. Army [DLA^{+}16]. Arrays [Grö20, HH17]. Art [EGB^{+}19, FLS^{+}20, HC16]. Ask [PDR^{+}17]. Assessing [ME14, VZR^{+}18]. Assessments [NMA^{+}18]. Assimilation [JCL^{+}20]. Assisted [Gri18]. Assurance [WL20]. Atmospheric [Juc14]. Atomic [OPA^{+}14]. Atomistic [BCOP^{+}18]. Attribute [TLR21]. Attribution [Kat14]. ATUS [Kim17]. ATUS-Commuting [Kim17]. autoAPI [Moo21]. Automated [SMS^{+}23, TLR21]. Automatic [MS14, TCD^{+}22]. Automating [FSBK18, JAGP14]. Avalanche [PDN18]. AvoPlot [Pet14].

BALTRAD [MHE^{+}18]. Barely [Bro14]. Based [BGB^{+}19, FLS^{+}20, Gen20, Han16, HKGvS21, KFV20, MA13, Moo21, OKVK17, RBB^{+}19, RLO18, RG21, SC20, Sta21, SMG19, ST15, BR21, BW16, FPF14, Gov16, GA13, JAGP14, OL17, Pet14, SBS^{+}18]. Basin [SA20, WLA19]. Basis [Sar17]. Batch [OKVK17]. Battery [GPKL^{+}20, ST15, SMT^{+}21]. Battery-Testers [GPKL^{+}20].

BayesFactorFMRI [Han21]. BayesFit [SSB19]. Bayesian [DF21, ES21, Han21, Kel20, LO16, OI16, Sch17, SSB19]. bayesint [DF21]. bayest [Kel20]. BayesTwin [Sch17]. bbsBayes [ES21]. BEAT [NMA^{+}18]. before [AS15]. Behavioral [Gri18, WH22]. Behaviours [Pas16]. Benthic [Kib16]. Best [SM14b]. Beta [DF21]. between [EA17, PKB19]. Big [TCH14]. Billie [Tau18]. Biodiversity [NMA^{+}18]. Biology [CB16b]. Biquandle [FW18]. Bird [ES21]. Bivalent [ME14]. Blind [VLG^{+}14]. Blood [DB17, KLC21, XXX16]. Boolean [OK20]. BoolSi [OK20]. Bottleneck [PFLG21]. Brain [RD22]. Breeding [ES21]. Brightness [SS18]. BST [ME14]. Build [PGR20]. Building [BCHR15, Tau18].

C [DAT^{+}21, LO16, MBA19, PDN18,

SGPHD^{+}17]. Cactus [LBAS14]. Calculate [Gen20, HKGvS21, Kim17]. Calculating [DF21]. Calculation [FW18, GOB16, TLR21]. Calculator [AV19]. Calibration [BGB^{+}19, LO16]. Caliko [LVSF16]. CALPHAD [OL17]. CALPHAD-based [OL17]. Cambridge [EBS17]. Camera [BR21]. Cameras [Wag17]. Capture [SZP15, SGPHD^{+}17]. Cardiovascular [DS16]. Carlo [SBV^{+}15, PDN18]. CARMA [Gir14]. Case [SC20]. Case-Based [SC20]. Cells [CNS^{+}21]. Cellular [KLC21]. CERF [VZR^{+}18]. CESER [Fer21]. CFD [MBLA19]. Challenges [ZZF^{+}14]. Change [CLW19, FS19, KFV20, PGR20, VLC^{+}18]. channel [SGPHD^{+}17]. Channeling [CB16a]. Characterization [Mar16]. Charting [BW16]. Chemical [dBHD17]. Children [ME14]. CHLOE [MS14]. Chromatography [Meu16]. CIT [Luk21]. Citation [Kat14]. Classification [BHF16, ZO16]. Classifiers [HR17]. Classifying [Kib16]. Client [HG23].

Client/Server [HG23]. Climate [Daw16a, HKGvS21]. Climatological [PdV21]. Cluster [Fer21]. Clustering [LW20]. CMakeCatchTemplate [DAT^{+}21]. Coarse [dBHD17]. Coarse-grained [dBHD17]. Code [BC17, DV14, Kim17, TCH14, SMG19]. Codes [DBB^{+}14]. Coding [ASL13, Gri18, Pas16]. Codon [RSR23]. Coefficients [Law17]. Coherence [NS20]. Colibri [vGvdB16]. Collection [ST15]. collections [FPF14]. Collective [SMG19]. Colloids [dBHD17]. Comfort [HH20, HH20]. Commitment [CKDG20]. Commitment/Economic [CKDG20]. Common [MBA19]. Communication [RD22]. Community [BCHR15, CB16a, DLR^{+}15, SMC^{+}22, WKD^{+}19, ZZF^{+}14]. Commuting [Kim17, WS15, Kim17]. Comparing [LW20]. Comparison [Han21]. Compartmental [FSBK18]. Complex [FS19, LV13, SC20]. COMPLEX-IT [SC20]. Component [WL13]. Comprehensive [SBS^{+}18]. Compressive [OPA^{+}14]. Computation [Coe17, DB15, Wan19]. Computational [GPP^{+}14, OL17, SM14b, FSV^{+}13]. Computations [CE19, Daw16b, NHL^{+}22]. Compute [Fer21]. computed [NHL^{+}22]. Computer [Gri18, Coe13, WH22]. Computer-Assisted [Gri18]. Computerized [ME14]. Computing [RWE^{+}20]. Concatenating [Sch21]. Concealed [Luk21]. Concerns [Kat14]. Concurrent [GPP^{+}14, SPG^{+}15]. Conditions [LFD^{+}22]. Conducting [LAB^{+}14]. ConfBuster [BVL18]. Configuration [Ned17]. Conformational [BVL18]. ConIII [LD19]. Connected [WL13]. Constrained [CKDG20]. Consumption [LPG^{+}18]. Container [SMS^{+}23]. Context [LAB^{+}14]. Continuous [GPP^{+}14, Gir14, SPG^{+}15]. Contributions [CB16a, EA17]. Control [Gar19, HWX^{+}16, WAH^{+}16]. Convenient [LD19]. Converting [XXX16]. Coordinates [DB15]. Copy [VLPV19]. Core [vGvdB16]. Correction [Han21, IKKC16]. Corrections [PAT22]. Correlation [LS21]. Counting [AMG21, CNS^{+}21]. Cover [CLW19, KFV20, VLC^{+}18]. CowLog [Pas16]. Creating [AS15, MYM21]. Creation [Grö20]. Credible [DF21]. Credit [Kat14, KS15]. Cross [FS16, LS21, MYM21, NS20, Pas16]. Cross- [LS21]. Cross-Platform [FS16, NS20, Pas16]. Cross-Tabulations [MYM21]. CSVDataMerge [Sch21]. CUDA [AC17]. Culture [CHI17, FPF14]. CURSAT [Gen20]. Curves [Gen20, WVTTG21]. Custom [BW16].

D [DB17, DB15, ML16, WAH^{+}16]. Data [AL16, Arg15, AMB19, BLDC^{+}19, BW16, CZZ19, Daw16a, DVS20, ES21, Far16, FLS^{+}20, FSBK18, FS16, Gen20, GPKL^{+}20, GDP18, HW19, HC16, HFM^{+}21, HGS21, HWX^{+}16, HMB23, JCL^{+}20, Juc14, KWV^{+}20, Kib16, KHD^{+}16, LAB^{+}14, ML16, MRX14, Moo21, PGR20, PBF^{+}16, PL20, PM17, Rab20, RG21, RML17, Sch21, Sch17, SR20, SMG19, SZP15, Wag17, Wan19, WTLB19, WL20, ZO16, FSV^{+}13, JAGP14, PFLG21, SSB19, Han22]. Database [NHL^{+}22]. Databases [SM14a]. DataDeps.jl [WTLB19]. DataExplore [Far16]. Datasets [HH17, MS14, PdV21]. Decadal [IKOC14, IKKC16]. Decay [WVTTG21]. Decision [HGHR20, SMG19]. Decision-Making [SMG19]. Deconvolution [Rab20, WVTTG21]. Define [SMC^{+}22]. Defocusing [BR21]. Defocusing-based [BR21]. DefocusTracker [BR21]. Degrees [EBS17]. Demeter [VLC^{+}18]. Deployment [Luk21]. Depth [Kib16]. Depth-tagging [Kib16]. Design [VEV^{+}19]. Designs [CC20].

Detection [BBJ17, Han16, MS14, TLR21]. Determine [HB17]. Determining [CNS^{+}21, Law17]. Developed [MOM21]. Developing [EGB^{+}19]. Development [AC17, CCH^{+}14, CB16b, GPP^{+}14, PM17, SPG^{+}15, SBV^{+}15]. Difference [BCOP^{+}18]. Differential [RN17]. DifferentialEquations.jl [RN17]. Diffusion [CZZ19]. Digital [Haz14, Kat14, KLC21, TS14]. Dilemma [KCH^{+}16]. Dimensional [WKC19, Wan19, Han22]. DiracQ [WS15]. Direct [JCL^{+}20]. Directory [OHM14]. Disaggregation [VLC^{+}18]. Discover [EA17]. Discovery [SMS^{+}23, VHT^{+}19]. Discrete [BC17]. Discussing [KP14]. Disequilibrium [PKB19]. Disparity [Jay21]. Dispatch [CKDG20]. Distributed [OK20]. Distributions [DF21]. Divided [HS18]. DiViDu [HS18]. Do [LMOB^{+}22]. Documents [DPL17]. DoE.MIParray [Grö20]. Domain [NS20]. Dot [RLO18]. Downscaling [LVH^{+}18]. Driven [PBF^{+}16]. Driver [vGFvNvA19]. Drop [DM16]. Drosophila [NL13]. DSCImageCalc [Law17]. DST [RD22]. Dual [HS18]. Dye [Wag17]. Dynamic [Gov16, Rab20, RD22]. Dynamics [AMB19].

e-MERLIN [Arg15]. e-Research [KP14]. eAnalytics [Gov16]. Earth [BLDC^{+}19, Han22, KILE^{+}21, RG21]. EasyVVUQ [RWE^{+}20]. ecg [DS16]. ecg-kit [DS16]. Ecological [SA20]. Economic [CKDG20]. Ecosystem [RN17, DPL17]. Ecosystems [Sto16]. EDI [PBF^{+}16]. Editor [Haz14, PBF^{+}16]. Education [Far16]. Effect [Kel20, SZH19]. Effect-Size [Kel20]. Effects [EA17, Haz14, TCH14]. Efficiency [CNS^{+}21]. Efficient [vGvdB16, GRZ13]. Efficiently [PL20]. Efflux [FSBK18]. Efforts [TCH14]. EIT-MESHER [DAF^{+}20]. Electrical [BS22]. Electron [GDP18]. Electrophysiology [Wag17]. Elephant [VLG^{+}14]. Embarrassingly [HG23]. Embedding [MOM21]. Embo [PFLG21]. Empirical [VLG^{+}14, PFLG21]. Enabling [GDP18]. encoding [SGPHD^{+}17]. Encyclopedia [SMC^{+}22]. End [BZSH21]. End-to-End [BZSH21]. Endpoint [Moo21]. Energy [Gov16, HGS21, VZR^{+}18, GDP18]. Engineering [DV14, HH15, LMOB^{+}22, GA13]. enhanced [TS21]. Enhancer [Haz14]. Ensembles [Kra16]. Entropy [LD19]. Environments [SM14b]. EOF [Daw16a]. eofs [Daw16a]. Equations [RN17]. Errors [Fer21]. Essence [GA13]. Estimated [Fer21]. Estimation [SZH19]. esy [PL20]. esy-osmfilter [PL20]. Evaluation [IKOC14, IKKC16, KILE^{+}21, VLG^{+}14]. Event [TLR21]. Executables [SMS^{+}23]. Expansion [VZR^{+}18]. Experience [CB16a]. Experiences [DV14, EGB^{+}19, KCL^{+}14, KCWD^{+}16, KCN^{+}16, KNG^{+}18, KDH^{+}19, PGR20, SBV^{+}15, ZZF^{+}14]. Experimental [AL16, CC20, Sch21]. Experiments [AMB19, HS18, FSV^{+}13]. Explorative [HLR15]. Exploratory [HGHR20]. Explorer [MA13]. Exploring [LW20, RML17]. ExpressInHost [RSR23]. Expression [RSR23]. ExpTimer [LV13]. Extendable [CSR^{+}23]. Extensible [MRX14, SM14b, Pet14]. External [dBHD17]. Extract [PL20]. Extraction [DPL17, Liu17]. Eye [Jon18]. Eye-Tracker [Jon18]. Eyex [Jon18].

FABRIK [LVSF16]. facilitate [LV13]. Fan [RDBC23]. Fan-Slicer [RDBC23]. Fast [RDBC23, vGFvNvA19, MBA19]. FastGAPP [RG21]. Feasibility [VZR^{+}18]. Feature [RN17, MA13]. Feature-Rich [RN17]. Features [AMG21, EA17]. FEM [DAF^{+}20]. Fermat [BG15]. Fidimag [BCOP^{+}18]. Field [Daw16b, HS18, WKD^{+}19]. Fields [Liu17, WKC19, dBHD17]. Figure [CC20]. File [Ned17]. Files [CFC^{+}17, Sch21]. Filter [BHF16]. Finding [BG15]. Finite [BCOP^{+}18]. Fire [AV19]. First [KCL^{+}14]. Flagging [HWX^{+}16]. Flexgram [vGvdB16]. Flexible [PGR20]. Flight [EBS17]. Flow [DB17, KLC21, Liu17, LS21, Meu16, MOM21, RDB21]. Flowtracks [ML16]. Fluid [AMB19, Meu16]. FluidDyn [AMB19]. FluidFFT [MBA19]. FluidSim [MBLA19]. fMRI [Han21, WL20]. Force [OPA^{+}14]. Forecasting [HKGvS21, MR21]. Forecasts [Kra16]. Forests [OKVK17]. FOrmatted [PAT22]. Fortran [AC17]. Forward [PMM15]. Fourier [MBA19, DB15, NS20]. Fourier-Domain [NS20]. Fourteen [Bro14]. Fourth [KNG^{+}18]. Framework [AMB19, FLS^{+}20, GPP^{+}14, HG23, Han22, HR17, JCL^{+}20, KILE^{+}21, KP14, KCH^{+}16, PKB19, SBS^{+}18, SPG^{+}15, SMC^{+}22, Tau18, UR19, VLG^{+}14, VHT^{+}19, VBAF16]. Frameworks [DV14]. Free [FS16, KILE^{+}21, Sch21, WAH^{+}16]. Freedom [EBS17]. Freshwater [SA20]. Freva [KILE^{+}21]. friendly [TS14]. Full [SS18]. Function [Sar17, TCD^{+}22]. Future [DBB^{+}14, VZR^{+}18].

Gage [RML17]. GAIL [TCD^{+}22]. Ganglion [CNS^{+}21]. Gas [Meu16]. GCAM [BLDC^{+}19]. gcamdata [BLDC^{+}19]. gcamland [CLW19]. Geiger [PDN18]. Genefer [BG15]. General [Far16]. Generalized [BG15, JNR17]. Generate [DVS20, Gen20]. Generating [Kra16]. Generation [DAF^{+}20]. Generator [Ned17, Jay21]. Generic [MRX14, MYM21, Gil17]. GENESIS [Gil17]. Genome [OI16]. Genome-Wide [OI16]. Geochemical [RG21]. Geospatial [KHD^{+}16, VZR^{+}18]. Geotagging [Kib16]. GIFT [SGPHD^{+}17]. GIFT-Grab [SGPHD^{+}17]. Giles [DPL17]. GIS [Kib16]. GitHub [SPG^{+}15]. Global [Daw16b, DVS20, LVH^{+}17, LVH^{+}18, VHT^{+}19]. Glyph [QGA19]. GNU [Rab20]. Go [BZSH21]. Grab [SGPHD^{+}17]. grained [dBHD17]. gram [vGvdB16]. Graph [RD22]. Graphic [Haz14]. Graphical [SM20]. Grid [HB17, SR20]. Gridsampler [HB17]. Guaranteed [TCD^{+}22]. Guided [Rab20]. GuiTeNet [SM20].

Habitat [Kib16]. Hand [FHRS21].

Hand-Over [FHRS21]. HANDE [SBV^{+}15]. Handling [RDB21]. Hankel [BC17]. Harmonics [Daw16b]. Harmonize [KWV^{+}20]. Harnessing [VEV^{+}19]. Heart [CRL21, vGFvNvA19]. Heat [VEV^{+}19]. HELM [MOM21]. HELMpy [MOM21]. Heterogeneous [PdV21]. Hierarchical [ES21]. High [Daw16b, GPKL^{+}20, Mat18, MBLA19, RWE^{+}20, Raß20]. High-Level [Daw16b, Mat18]. High-Performance [MBLA19]. High-Precision [GPKL^{+}20]. Hindcast [IKOC14, IKKC16]. HiPPSO [Raß20]. Historical [DVS20, OHM14]. Holomorphic [MOM21]. Homogeneity [HMB23]. Homology [FW18]. Host [RSR23]. HPC [CCH^{+}14, DBB^{+}14, MBA19]. HTML5 [MA13]. Hub [WKD^{+}19]. Human [BLDC^{+}19, Sto16, WH22]. Human-computer [WH22]. HydroCloud [RML17]. Hydrodynamics [dBHD17]. Hydrologic [LVH^{+}17, VHT^{+}19]. Hydropower [WLA19].

Ideal [LFD^{+}22]. IJBlob [WL13]. Image [DM16, Haz14, HR17, Law17, LS21, MS14, Mar16, MGMP16, OKVK17, TS14, TS21, Haz14]. ImageJ [CNS^{+}21, DM16, HH15, OKVK17, WL13, ZO16]. Images [Cre22, Han21, Liu17, OPA^{+}14]. ImageSURF [OKVK17]. Imaging [FS16, Wag17]. Impacts [Sto16]. Impaired [EGB^{+}19]. Implementation [LVSF16, NS20, NHL^{+}22, RD22]. Implemented [BHF16]. Implementing [Han21]. Including [MOM21]. Index [HKGvS21, LPG^{+}18]. Indicators [SA20]. Industrial [Gar19]. Industry [Gov16]. Inference [CRL^{+}19, OI16, Sch17, SSB19]. Information [Luk21, PFLG21, Tau18]. Infrastructure [CHI17, KP14, SM14b, UR19]. Input [BLDC^{+}19, PDR^{+}17, Sta21]. Input-Output [Sta21]. Inputlog [VLPV19]. Inputs [DVS20]. Inquiry [SC20]. Insert [HW19]. Integrated [BLDC^{+}19, HS18, NMA^{+}18, PM17]. Integrating [LS21, SA20]. Integration [GPP^{+}14, SPG^{+}15, TCD^{+}22, TCH14, UR19]. Interaction [WH22]. Interactive [BGB^{+}19, BW16, PdV21, RBB^{+}19, WH22]. Interface [Jon18, LD19, Mat18, SS18, SM20]. Interference [ME14]. Interpolator [WKC19]. Interpreting [RG21]. Interrelated [EA17]. Intervals [DF21]. Introducing [HGS21]. Introduction [KILE^{+}21]. Intuitive [Hug14]. Inverse [LVSF16, LD19]. iOS [LPG^{+}18]. Ising [LD19, LD19]. Ising-Type [LD19]. Issues [LBAS14]. Item [Sch17]. Item-Level [Sch17]. Iterated [KCH^{+}16].

Janus [KFV20]. Japanese [PDR^{+}17]. Java [ST15]. Java-Based [ST15]. Job [GRSLL16]. JSON [KS15, Moo21]. JSON-LD [KS15]. Jug [Coe17]. Julia [CE19, CSR^{+}23, JNR17, RN17]. Julian [HKGvS21]. Julian-Oscillation [HKGvS21]. Just [Bro14].

KAMG [XXX16]. Kinematics [LVSF16]. Kinematogram [RLO18]. kit [DS16]. Knife [DLA^{+}16]. Know [LMOB^{+}22]. Komadu [SZP15]. KPP [AC17].

L1C [SS18]. Lab [CB16b]. labeled [HH17]. Laboratory [AV19]. Land [CLW19, DVS20, KFV20, VLC^{+}18]. Lane [vGFvNvA19]. Langevin [RLWP16]. Language [CSR^{+}23, FS19, HC16]. Large [BG15, CKDG20, DV14, ZZF^{+}14]. Large-Scale [CKDG20]. LD [KS15]. Leaping [AS15]. Learning [HR17, KDH^{+}19]. LEED [GDP18]. LEEM [GDP18]. Lessons [BCHR15]. Level [Daw16b, Han21, Jay21, Mat18, Sch17]. Libraries [MBA19]. Library [BW16, CZZ19, Daw16a, Daw16b, Gar19, HGHR20, HC16, HFM^{+}21, JLZ21, LVSF16, PL20, PM17, RWE^{+}20, TCD^{+}22, TLR21, WL13, GRZ13, JAGP14]. Lifecycle [LAB^{+}14]. light [GRZ13]. light-weight [GRZ13]. Lightweight [GRSLL16]. LIIS [FPF14]. Linear [PDN18]. Linkage [PKB19]. Linkages [KWV^{+}20]. Liquid [Meu16]. Listen [BZSH21]. Listening [BZSH21, SBS^{+}18]. Load [MOM21]. Locking [RD22]. Longbow [GRSLL16]. Looking [AS15]. Low [GDP18]. Low-energy [GDP18].

Machina.NET [Gar19]. Macro [MYM21]. Macrocycle [BVL18]. Macros [FCY20]. Madagascar [FSV^{+}13]. Madden [HKGvS21]. Magni [OPA^{+}14]. Mahotas [Coe13]. Making [HGHR20, SMG19]. Management [LAB^{+}14, ML16, MRX14, SA20, WLA19, JAGP14]. Manipulation [RDB21, WS15]. Manual [HYG16]. Many [HGHR20]. Many-Objective [HGHR20]. Maple [DB15]. Mapping [Kib16, Sto16]. Marine [Sto16]. Markov [RLWP16]. Massive [HG23]. MatCal [LO16]. Materials [VEV^{+}19]. MATH [GOB16]. Mathematical [SdSS16, SMT^{+}21]. MATLAB [Jon18, RG21, RD22, SM14a, TS14, TS21, TCD^{+}22, BC17, DLA^{+}16, DS16, LO16, Mar16, Sar17]. MATLAB(R) [Rab20, SS18]. matplotlib [Pet14].

Matrices [XXX16]. Matrix [ZO16]. Maximum [LD19]. Means [Kat14]. Measure [AV19, DM16]. Measuring [VLPV19]. MECCA [AC17]. Media [Gir14]. Mediation [FCY20, YL17]. Mediators [YL17]. MEDINA [AC17]. Memory [ST15]. Men [VLG^{+}14]. Mentorship [TCH14]. MERLIN [Arg15]. Mesh [DAF^{+}20]. MESHER [DAF^{+}20]. message [GRZ13]. Meta [Han21]. Meta-Analysis [Han21]. Metadata [PBF^{+}16]. Meteorological [Daw16a, PdV21, Wan19]. Method [LS21, MOM21]. Methodologies [DBB^{+}14]. Methods [GOB16]. Metis [KWV^{+}20]. Micro [NHL^{+}22]. Micro-computed [NHL^{+}22]. Micromagnetic [BCOP^{+}18, VBAF16]. Micromagnetics [VBAF16]. Microorganisms [RSR23]. Microscopy [MS14, OPA^{+}14, ZO16]. MINRES [Cho14]. MINRES-QLP [Cho14]. MiToBo [MGMP16]. mixchar [WVTTG21]. ML [PDR^{+}17]. ML-Ask [PDR^{+}17]. mma [YL17]. Mode [PDN18]. Model [BLDC^{+}19, CKDG20, EML21, HH20, LVH^{+}17, LFD^{+}22, Rab20, Sto16, Tau18, VZR^{+}18, VLC^{+}18, WLA19]. model4you [SZH19]. Modeling [DVS20, HGHR20, HGS21, KILE^{+}21, KFV20, RLWP16, GA13, SSB19]. Modelling [CLW19, FS19, Gil17, MR21, PDN18, SC20, SMT^{+}21, vGvdB16]. Models [LD19, RBB^{+}19, UR19]. Modular [BR21, HR17, HFM^{+}21, HYG16, MBLA19]. Moirai [DVS20]. Moisture [BGB^{+}19]. MongoDB [HW19]. Monitoring [LPG^{+}18, MA13]. Monte [SBV^{+}15, PDN18]. MOOSE [SPG^{+}15]. Moosh [DLA^{+}16]. Motion [WAH^{+}16]. Motor [RD22, VLPV19]. MPWide [GRZ13]. MUFOS [PAT22]. Multi [KWV^{+}20, Sta21, Wan19, LV13, SGPHD^{+}17]. multi-channel [SGPHD^{+}17]. Multi-Dimensional [Wan19].

Multi-Regional [Sta21]. Multi-Sectoral [KWV^{+}20]. multi-step [LV13]. Multidimensional [SR20, FSV^{+}13]. Multilayers [DLA^{+}16]. Multilingual [VLPV19]. Multiphysics [DV14, HFM^{+}21]. Multiple [FCY20, Han21, Wag17, YL17, PAT22]. Multiprocessing [Han21]. Multivariate [SR20]. MurCSS [IKOC14, IKKC16]. MWA [HYG16]. Myex [Jon18].

nanoHUB.org [ZZF^{+}14]. Nansat [KHD^{+}16]. Navigating [CFC^{+}17]. nd [Han22]. Network [CKDG20, HW19, HR17, LFD^{+}22, RD22]. Network-Constrained [CKDG20]. Networking [MHE^{+}18]. Networks [EA17, OK20, SM20, GRZ13]. Neural [EA17, HR17]. NeuroCharter [EA17]. NMME [Kra16]. Noisy [CRL^{+}19, vGFvNvA19]. NOMA [LFD^{+}22]. Non [LFD^{+}22, WS15]. Non-Commuting [WS15]. Non-Ideal [LFD^{+}22]. North [ES21]. Novelty [MS14]. Novonix [GPKL^{+}20]. npFEM [KLC21]. Numerical [DLA^{+}16, GOB16]. Numerous [Haz14].

OACoder [ASL13]. Object [MBLA19]. Object-Oriented [MBLA19]. Objective [HGHR20]. Observation [Gri18, Han22]. ObsPy [TLR21]. Ocean [RBB^{+}19, vdB19]. Oceanographic [Daw16a, RBB^{+}19]. OCR [DPL17]. OCT [NS20]. Octave [DLA^{+}16, Rab20, SS18, SM14a]. Octave/Matlab [DLA^{+}16]. Octave/MATLAB(R) [SS18]. oemof.tabular [HGS21]. Off [vGFvNvA19]. Off-the-Shelf [vGFvNvA19]. Office [OHM14]. oflibnumpy [RDB21]. oflibpytorch [RDB21]. Older [RD22]. OLR [HKGvS21]. OLR-Based [HKGvS21]. OMI [HKGvS21]. One [CC20]. Online [BZSH21, JAGP14]. Open [AMG21, AMB19, BVL18, Coe13, EML21, Gri18, HS18, HYG16, KCH^{+}16, LPG^{+}18, Liu17, LS21, LO16, MOM21, PAT22, PdV21, PDR^{+}17, SA20, SM14a, SBV^{+}15, Sto16, TCD^{+}22, TCH14, Wan19, WAH^{+}16, FSV^{+}13]. Open-Source [AMG21, AMB19, BVL18, Gri18, HYG16, PAT22, SBV^{+}15, TCD^{+}22, SM14a, WAH^{+}16, FSV^{+}13]. OpenOpticalFlow [Liu17]. OpenOpticalFlow PIV [LS21]. OpenStreetMap [PL20]. Optical [LS21, NS20, RDB21]. Optics [DLA^{+}16]. Optimization [EML21, Meu16, Raß20, SdSS16, TCD^{+}22]. Optimizer [EML21]. OPTN [HW19]. Organ [HW19]. Orientated [KHD^{+}16]. Oriented [MBLA19]. Orthogonal [Grö20]. Oscillation [HKGvS21]. osmfilter [PL20]. Outlier [HWX^{+}16]. OutlierFlag [HWX^{+}16]. Output [FLMR16b, FSBK18, Sta21].

Pack [Cho14]. Package [AL16, BCOP^{+}18, BLDC^{+}19, CLW19, CRL21, DB17, DF21, ES21, Fer21, GPKL^{+}20, Grö20, HKGvS21, HMB23, KFV20, Kel20, KHD^{+}16, Kra16, LD19, LVH^{+}18, ML16, MBLA19, MOM21, OPA^{+}14, RDBC23, RLWP16, Sch17, SZH19, SdSS16, VEV^{+}19, WL20, WVTTG21, WS15, YL17, vdB19, PFLG21]. Packages [HGS21]. Palabos [KLC21]. Palabos-npFEM [KLC21]. Paradigm [HS18]. Parallel [Coe17, HG23, NHL^{+}22]. Parameter [EML21]. Parameters [Ned17]. ParaView [Juc14]. Parchar [Mar16]. Parser [OHM14, OHM14]. Particle [BR21, Hug14, LS21, ML16, Raß20, TS14, TS21]. Particles [Mar16]. Partitioning [PKB19]. passing [GRZ13]. Patching [PMM15]. PBTK [EML21]. Pendent [DM16]. Pendent Drop [DM16]. Perception [HH20]. PerfAndPubTools [FLMR16a]. Performance [FLMR16a, MBLA19, RWE^{+}20, SdSS16]. Performant [RN17]. Perprof [SdSS16].

Perprof-py [SdSS16]. Personalised [SZH19]. Perspective [BL14]. Petrological [RG21]. PFHub [WKD^{+}19]. Phase [RD22, WKD^{+}19]. Phase-Field [WKD^{+}19]. Phase-Locking [RD22]. Photo [Kib16, NL13]. Photo-transduction [NL13]. Photodiodes [PDN18]. Photos [Kib16]. Physiology [vGFvNvA19]. PhysioNet [SM14a]. Picasso [HR17]. PINTS [CRL^{+}19]. Pipeline [Arg15, RD22]. PIVlab [TS14, TS21]. Pixel [OKVK17]. Pixel-Based [OKVK17]. Planes [RDBC23]. Platform [BZSH21, FSBK18, FS16, NS20, Pas16, SC20]. Please [EGB^{+}19, GDP18]. Plots.jl [CSR^{+}23]. Plotting [CFC^{+}17, CSR^{+}23, Pet14]. Plugin [DM16, HH15, OKVK17, ZO16]. Plugins [CNS^{+}21]. POD [OHM14]. Polar [DB15]. Polarization [SS18]. Populations [PKB19]. Porous [HH15]. Portage [HFM^{+}21]. Post [OHM14]. Postcode [ASL13]. Posterior [RD22]. Posterior-to- [RD22]. Power [BHS18, CKDG20, MOM21]. PowNet [CKDG20]. Practice [KCL^{+}14, KCWD^{+}16, KCN^{+}16, KNG^{+}18, KDH^{+}19]. Practices [SM14b]. Pre [AC17]. Pre-processor [AC17]. Precision [GPKL^{+}20, Raß20]. preconfig [Ned17]. Predictive [Kra16]. Preparation [BLDC^{+}19]. Preparing [AL16, TLR21, WH22]. prepdat [AL16]. Preprocess [GPKL^{+}20]. Preprocessing [CRL21]. Preserving [WH22]. Primes [BG15]. Prisoner [KCH^{+}16]. Probabilistic [Kra16, EBFS17, CRL^{+}19]. Probable [BG15]. Problems [DB17]. procedures [LV13]. Process [HR17, PGR20]. Processes [RLWP16, CB16b]. Processing [AMB19, DS16, DVS20, FLS^{+}20, Hug14, KHD^{+}16, Mar16, MGMP16, SS18, SM14a, SR20, WAH^{+}16, SGPHD^{+}17]. processor [AC17]. Procurement [HW19]. Produced [GPKL^{+}20]. Producing [Kib16]. Production [VZR^{+}18]. Products [Kat14].

Profile [SdSS16]. Program [Gen20, Liu17, LS21, RG21, Sch21]. Programmable [Mat18]. Programming [CSR^{+}23, Gar19, HC16]. Programs [BG15]. Project [BCHR15, DAT^{+}21, FSV^{+}13, OHM14, SBV^{+}15, vGFvNvA19]. Propagating [AV19]. Proteins [RSR23]. Prototypical [Tau18]. Provenance [SZP15]. Pseudoreplicates [Gen20]. psychophysical [SSB19]. PsychoPy [Han16]. PsychoPy-Based [Han16]. Publication [MYM21, JAGP14]. Publication-Quality [MYM21]. Publishing [FLMR16a]. Pulling [PMM15]. Py [FLS^{+}20, HC16, SdSS16]. Py-ART [FLS^{+}20, HC16]. PyBaMM [SMT^{+}21]. pycalphad [OL17]. Pycuda [RDBC23]. PyDDA [JCL^{+}20]. pydiffusion [CZZ19]. pyfMRIqc [WL20]. pyHomogeneity [HMB23]. Pymrio [Sta21]. pyObs [Gri18]. pyparty [Hug14]. PyPSA [BHS18]. Pyrad [FLS^{+}20]. PyRDM [JAGP14]. Python [AMB19, BHF16, BHS18, CRL21, CZZ19, Coe17, DB17, DF21, EML21, GPKL^{+}20, GDP18, HGHR20, HW19, HC16, HKGvS21, HH17, Hug14, HLR15, HMB23, JAGP14, KFV20, KHD^{+}16, LD19, LVH^{+}18, MBA19, MBLA19, MOM21, OL17, OPA^{+}14, PFLG21, PL20, RDB21, SGPHD^{+}17, SdSS16, Sta21, SMT^{+}21, VEV^{+}19]. Python-based [JAGP14]. Pythonic [JCL^{+}20]. PyVDT [Han16].

QLP [Cho14]. qtl [Bro14]. Qualitative [SMG19]. Quality [HWX^{+}16, MYM21, WL20]. Quandle [FW18]. Quantification [RWE^{+}20]. Quantum [SBV^{+}15, WS15].

R [AL16, BLDC^{+}19, Bro14, CLW19, ES21, FW18, Fer21, FCY20, Grö20, Kel20, PKB19, RLWP16, Sch17, SZH19, WVTTG21, YL17]. R/qtl [Bro14]. Radar [FLS^{+}20, HC16, MHE^{+}18]. Radial [Sar17].

Random [OKVK17, RLO18]. Random-Dot [RLO18]. Rapid [GDP18, SS18]. Rarefaction [Gen20]. Rate [AV19, CRL21, vGFvNvA19]. Rates [Meu16]. Rating [Gir14]. Ratios [DF21]. Raw [WL20]. Reactions [dBHD17]. Ready [Luk21]. Ready-to-Use [Luk21]. Real [FLS^{+}20, Gar19, NS20, SGPHD^{+}17, vGFvNvA19]. Real-Time [FLS^{+}20, Gar19, NS20, vGFvNvA19, SGPHD^{+}17]. Recombinant [RSR23]. Recommendations [DLR^{+}15]. Reconstruction [OPA^{+}14]. Recordings [BBJ17]. Reduction [Arg15, MR21]. Reference [NHL^{+}22]. Refinement [DAF^{+}20]. Regional [Sta21]. Register [Jay21]. Registry [AS15]. Regression [BHF16, OI16, QGA19]. Related [ML16]. Reliable [Cho14]. Remap [HFM^{+}21]. Remote [Cre22, GRSLL16]. Repeatable [WTLB19]. Repertory [HB17]. Replay [TDX^{+}20]. Report [KCWD^{+}16, KCN^{+}16]. Reporting [LPG^{+}18]. Representation [SR20]. Reproducibility [CHI17]. Reproducible [Cho14, Coe17, HGS21, KCH^{+}16, SM14b, VBAF16, WTLB19, FSV^{+}13]. Required [HB17]. Resampling [Gen20]. Resampling-Based [Gen20]. Research [AMB19, Cho14, CHI17, Far16, FHRS21, KP14, KDH^{+}19, LMOB^{+}22, PGR20, PBF^{+}16, RLO18, SMC^{+}22, SM14b, SMC^{+}22]. Reservoir [WLA19]. Resistivity [BS22]. Reslicing [RDBC23]. Resolution [Cre22]. Resolved [ZO16]. Response [Luk21]. Results [FLMR16a]. Retinal [CNS^{+}21]. Retrievals [JCL^{+}20]. Return [Moo21]. Reveals [RD22]. RGC [CNS^{+}21]. Rho [BS22]. Rhodium [HGHR20]. Rich [RN17]. River [WLA19]. RMPCDMD [dBHD17]. Robotic [JLZ21]. Robots [Gar19]. Robust [HGHR20]. Rocket [EBS17]. Rocketry [EBS17]. rOpenSci [BCHR15]. ROS [AV19]. RWebData [Mat18].

Sample [HB17, Kel20, FPF14]. Sampling [JNR17, OPA^{+}14]. Sandbox [FS19]. Sandboxing [WH22]. SAS [FCY20, MYM21]. Scaffolds [HH15]. Scalar [WKC19]. Scale [CKDG20, DV14]. Scales [DVS20, KWV^{+}20]. ScatterJn [ZO16]. Scatterplot [ZO16]. Scatterplot-Matrix [ZO16]. Scenario [SC20]. Schemes [BHF16]. Science [HKGvS21, KCL^{+}14, KCWD^{+}16, KCN^{+}16, KNG^{+}18, KDH^{+}19, LAB^{+}14, SM14b, WTLB19]. Scientific [BL14, Cho14, CB16a, CB16b, DLR^{+}15, GOB16, HWX^{+}16, Juc14, SBV^{+}15, SZP15, TCH14, VHT^{+}19, ZZF^{+}14, JAGP14, Pet14]. Scientist [BL14, KHD^{+}16]. Scientist-Orientated [KHD^{+}16]. Scientists [RG21]. Scikit [HLR15]. Scikit-spectra [HLR15]. Scraping [WH22]. Screening [WLA19]. scriptable [Coe13]. Scripts [HW19]. Search [BVL18]. Second [Han21, KCWD^{+}16]. Second-Level [Han21]. Sectoral [KWV^{+}20]. Sediment [WLA19]. Sedimentological [RG21]. SedSim [WLA19]. SeFo [Kra16]. Segmentation [OKVK17]. Segmentations [Law17]. Segmented [DAF^{+}20]. Seismic [TLR21]. Self [LPG^{+}18]. Self-Reporting [LPG^{+}18]. SEMAT [GA13]. Sensing [Cre22]. sensitive [Wag17]. Sensors [BGB^{+}19, vGFvNvA19]. Sequence [Han16]. Series [CRL^{+}19, FLMR16b, HMB23]. Server [HG23, PdV21]. Setting [AV19]. Setup [WTLB19]. Seven [ST15]. Shape [ME14, WL13]. Shaped [RDBC23]. Shapefiles [Kib16]. Shelf [vGFvNvA19]. Shift [RD22]. Short [TCH14]. Short-term [TCH14]. Signal [CRL21, DS16]. Similarity [Law17]. SimOutUtils [FLMR16b]. Simple [Gen20, PDN18, Sch21, CNS^{+}21]. Simplifying [CCH^{+}14]. Simulation [BCOP^{+}18, CZZ19, DBB^{+}14, FLMR16b, HB17, KLC21, LBAS14, SC20, UR19, VBAF16, WLA19, Gil17]. Simulations [AMB19, MBLA19, OK20, dBHD17]. Simulator [EBS17, NL13, PDN18, TDX^{+}20, HH20]. Single [BR21]. Single-Camera [BR21]. Sites [WH22]. Six [EBS17]. Six-Degrees-of-Freedom [EBS17]. Size [HB17, Kel20]. Skills [VLPV19]. Skipgram [vGvdB16]. Slicer [RDBC23]. Small [CB16b]. SmaRT [Luk21]. SmaRT-CIT [Luk21]. Smartphone [Luk21]. SMOS [SS18]. Smurf [MR21]. SnappySonic [TDX^{+}20]. Social [Kat14, SC20, SA20]. Social-Ecological [SA20]. Software [AS15, AMG21, BL14, BCHR15, Cho14, Coe17, CCH^{+}14, CB16a, CB16b, DLR^{+}15, DV14, DBB^{+}14, EA17, FLMR16a, FHRS21, Gir14, Gri18, HH20, KCL^{+}14, KCWD^{+}16, KCN^{+}16, KNG^{+}18, KDH^{+}19, Kib16, KLC21, LMOB^{+}22, LVSF16, Law17, LPG^{+}18, LAB^{+}14, LBAS14, MS14, ML16, MRX14, ME14, OHM14, PAT22, PDR^{+}17, SdSS16, SMC^{+}22, SBV^{+}15, SMG19, Sto16, SM14b, TCH14, UR19, VLG^{+}14, Wan19, WAH^{+}16, WL20, ZZF^{+}14, vGFvNvA19, Coe13, FSV^{+}13, GA13, JAGP14, LV13]. Soil [BGB^{+}19]. Solution [HS18]. Solve [DB17]. Solvers [MOM21]. Solving [LD19, RN17]. SomaLogic [CFC^{+}17]. Source [AMG21, AMB19, BVL18, EML21, Gri18, HS18, HYG16, LPG^{+}18, Liu17, LS21, LO16, MOM21, Moo21, PAT22, PDR^{+}17, SA20, SBV^{+}15, Sto16, TCD^{+}22, TCH14, Wan19, AC17, Coe13, FSV^{+}13, SM14a, WAH^{+}16]. source-to-source [AC17]. Spacing [NHL^{+}22]. Spatial [KWV^{+}20, LVH^{+}18, SR20]. Spatially [ZO16]. Specification [Moo21]. spectra [HLR15]. Spectram [Rab20]. Spectroscopic [FS16, Rab20]. Spectroscopy [HLR15]. Spherical [Daw16b]. Spline [WKC19]. Spread [AV19]. Sprint [CB16a]. SR4RS [Cre22]. Standard [Fer21]. Standardized [IKOC14, IKKC16]. Standards [PdV21]. Stata [Kim17]. State [Gil17, KDH^{+}19]. Stations [RBB^{+}19]. Statistical [AL16]. Stemming [Kat14]. step [LV13]. Stereo [Jay21]. Stimulus [RD22]. Stochastic [EBS17]. Storage [DPL17]. Stream [RML17]. Structuring [SMG19]. Studies [HB17, WH22]. Study [KCH^{+}16]. StudySandboxx [WH22]. Subjects [RD22]. Submission [GRSLL16]. Suite [Wan19, GDP18]. Summary [KCL^{+}14]. Super [Cre22]. Supercritical [Meu16]. Supportable [Cho14]. Supporting [HG23]. Supports [RG21]. Suppression [ME14]. Surface [DM16]. Survey [ES21, Kib16, Kim17]. Suspended [Mar16]. Sustainability [CCH^{+}14, KP14, VLG^{+}14, ZZF^{+}14]. Sustainable [BL14, Bro14, CB16b, DLR^{+}15, FHRS21, KCL^{+}14, KCWD^{+}16, KCN^{+}16, KNG^{+}18, KDH^{+}19, LBAS14]. svy freqs [MYM21]. Swarm [Raß20]. SWCalibrateR [BGB^{+}19]. Swiss [DLA^{+}16]. Symbolic [QGA19]. Synchronization [RD22]. Synthesis [BLDC^{+}19]. System [BHS18, DVS20, FS19, Gil17, HGS21, KILE^{+}21, MR21, SZP15, FPF14]. Systems [BLDC^{+}19, CKDG20, CHI17, IKOC14, IKKC16].

Tables [PAT22]. Tabulations [MYM21]. Tadarida [BBJ17]. tagging [Kib16]. Taking [vGFvNvA19]. Talk [CB16b]. Targeted [Kel20]. Task [Han16, HS18, ME14, VLPV19]. Taskfarm [HG23]. Tasks [ST15]. Teaching [AMB19]. Technological [Kat14]. Technology [VZR^{+}18]. Teetool [EBFS17]. Temperature [SS18]. Template [JLZ21, PBF^{+}16, DAT^{+}21]. Template-Driven [PBF^{+}16]. Temporal [LVH^{+}18]. Tension [DM16]. Tensor [CE19, SM20, vdB19]. Tensors.jl [CE19]. term [TCH14]. Test [BZSH21, HMB23, Luk21, ST15]. Testers [GPKL^{+}20]. Testing [ME14]. Tests [Kel20, PAT22, SBS^{+}18]. Tethys [LVH^{+}18]. Text [DPL17]. Textual [PDR^{+}17]. theory [GA13]. Thermal [WVTTG21]. Thermodynamics [OL17]. Thermoregulation [HH20]. Thickness [NHL^{+}22]. Third [KCN^{+}16]. Three [MYM21, WKC19]. Three-Dimensional [WKC19]. Three-Way [MYM21]. Ties [XXX16]. Time [CRL^{+}19, FLMR16b, FLS^{+}20, Gar19, HMB23, Kim17, Luk21, NS20, vGFvNvA19, SGPHD^{+}17]. timer [LV13]. Tissue [HH15]. Tobii [Jon18]. Tomography [NS20, NHL^{+}22]. Tool [AV19, ASL13, CFC^{+}17, CC20, Cre22, FS16, GRSLL16, GOB16, HB17, HYG16, HWX^{+}16, HH20, IKOC14, IKKC16, KWV^{+}20, MS14, MA13, Meu16, Moo21, NMA^{+}18, OK20, RBB^{+}19, RSR23, WH22, XXX16, EBFS17, GA13, Pet14, SSB19]. Toolbox [BR21, BBJ17, DS16, DB15, MGMP16, Rab20, RD22, SA20, SM14a, Sta21, Sar17]. Toolkit [HC16]. Tools [BVL18, FLMR16a, PGR20, QGA19]. Touch [EGB^{+}19]. Trabecular [NHL^{+}22]. Tracer [FSBK18]. Tracker [Jon18]. Tracking [BR21, BLDC^{+}19, ML16]. trajectory [EBFS17]. Transduction [CNS^{+}21, NL13]. Transfer [Jay21]. Transform [BC17, MBA19]. Transformation [RD22]. Transforms [DB15]. Transition [Rab20]. Transitions [Gil17]. Transitive [Kat14, KS15]. Transplant2Mongo [HW19]. Transplantation [HW19]. Treatment [SZH19]. Tricubic [WKC19]. Tuning [RSR23]. Twin [Sch17]. Two [Kel20]. Two-Sample [Kel20]. Type [LD19]. Typing [VLPV19].

UAV [LFD^{+}22]. UAV-NOMA [LFD^{+}22].

Uavnoma [LFD^{+}22]. Ultrasound [RDBC23, TDX^{+}20]. UN-CODE [SMG19]. Uncertainty [MR21, RWE^{+}20]. Unit [CKDG20]. Unsupervised [LW20]. Usage [MA13]. Use [CLW19, CCH^{+}14, KFV20, Kim17, Luk21, VLC^{+}18, WH22]. used [HH15]. User [CSR^{+}23, PGR20, SM20, TS14]. User-friendly [TS14]. Using [Daw16b, FCY20, OKVK17, PdV21, Wag17, vGFvNvA19, PFLG21, SSB19]. Utilities [FLMR16b].

v1.0 [CLW19]. vaCATE [FSBK18]. Validation [RBB^{+}19, RWE^{+}20]. Validator [RBB^{+}19]. VaMpy [DB17]. Variability [CRL21]. Variable [KWV^{+}20]. Variables [WS15]. Variational [OI16]. Various [DVS20]. Varying [Ned17]. Vector [WKC19]. Vectors [CNS^{+}21]. Velocimetry [LS21, ML16, TS14, TS21]. Velocity [Liu17]. ver [Gen20]. Verification [RWE^{+}20]. Versatile [Ned17]. Version [DVS20]. Vespucci [FS16]. via [Cho14, DB15]. Video [AMG21, Pas16, SGPHD^{+}17]. VIFECO [AMG21]. VIGoR [OI16]. Viral [CNS^{+}21]. Virtual [VBAF16]. Vision [Jay21, RLO18, Coe13]. Visual [Han16, HS18]. Visualisation [Juc14, Wan19]. Visualization [BW16, CC20, GOB16, Liu17, MA13, SZP15, Tau18]. Visualizing [HR17, SMG19]. Visually [EA17, EGB^{+}19]. Voltage [Wag17]. Voltage-sensitive [Wag17]. Vowel [FS19]. VScope [Wag17].

Walking [CB16b]. Water [LVH^{+}18, WLA19]. Waveforms [TLR21]. Way [MYM21]. Weather [FLS^{+}20, HC16, HKGvS21, MHE^{+}18]. web [FPF14, BGB^{+}19, BW16, CFC^{+}17, CC20, Gov16, GA13, Mat18, Moo21, RBB^{+}19, RLO18, RML17, SBS^{+}18, WH22]. web-based [FPF14, BGB^{+}19, Moo21, RBB^{+}19, RLO18, BW16, Gov16, GA13, SBS^{+}18]. Webcharts [BW16]. webMUSHRA [SBS^{+}18]. weight [GRZ13]. WEKA [BHF16]. WekaPyScript [BHF16]. Whisker [HYG16]. Wholemounts [CNS^{+}21]. Wide [OI16, GRZ13]. Wildfire [AV19]. Wind [Daw16b, JCL^{+}20]. Windspharm [Daw16b]. Withdrawals [LVH^{+}18]. Within [PKB19]. Workflows [HGS21]. Working [HC16, ST15]. Workloads [HG23]. Workshop [KCL^{+}14, KCWD^{+}16, KCN^{+}16, KNG^{+}18, KDH^{+}19]. WPTherml [VEV^{+}19]. Writing [VLPV19]. WSSSPE1 [KCL^{+}14]. WSSSPE2 [KCWD^{+}16]. WSSSPE3 [KCN^{+}16]. WSSSPE4 [KNG^{+}18]. WSSSPE5.1 [KDH^{+}19].

Xanthos [LVH^{+}17]. xarray [HH17]. XML [Moo21].

Years [Bro14].

References

Alvanos:2017:PMM

[AC17] Michail Alvanos and Theodoros Christoudias. MEDINA: MECCA development in accelerators — KPP Fortran to CUDA sourceto-source pre-processor. Journal of Open Research Software, 5(1):13–??, April 28, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                                     openresearchsoftware.metajnl.
                                                     com/articles/10.5334/jors.
                                                     158/.

Allon:2016:PPR

[AL16] Ayala Allon and Roy Luria. prepdat — an R package for preparing experimental data for statistical analysis. Journal of Open Research Software, 4(1):e43–??, November 25, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
134/.

Augier:2019:FPO

[AMB19] Pierre Augier, Ashwin Vishnu Mohanan, and Cyrille Bonamy. FluidDyn: a Python opensource framework for research and teaching in fluid dynamics by simulations, experiments and data processing. Journal of Open Research Software, 7(1):9–??, April 01, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
237/.

Apparicio:2021:VOS

[AMG21] Philippe Apparicio, David Maignan, and Jérémy Gelb. VIFECO: an open-source software for counting features on a video. Journal of Open Research Software, 9 (1):7–??, May 07, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
300/.

Argo:2015:PMD

[Arg15] Megan Argo. The e-MERLIN data reduction pipeline. Journal of Open Research Software, 3(1):e2–??, January 29, 2015. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        bp/.

Allen:2015:LBL

[AS15] Alice Allen and Judy Schmidt. Looking before leaping: Creating a software registry. Journal of Open Research Software, 3(1):e15–??, November 23, 2015. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        bv/.

Adnan:2013:POP

[ASL13] Muhammad Adnan, Alex Singleton, and Paul Longley. OACoder: Postcode coding tool. Journal of Open Research Software, 1(1):e5–??, October 08, 2013. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/511ba2c94d661/

.

Abouali:2019:FRC

[AV19] Abdelrahman Abouali and Domingos Xavier Viegas. Fire

                                        ROS Calculator: a tool to

measure the rate of spread of a propagating wildfire in a laboratory setting. Journal of Open Research Software, 7(1):24–??, July 29, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        221/.

Bas:2017:PTT

[BBJ17] Yves Bas, Didier Bas, and Jean-François Julien. Tadarida: A toolbox for animal detection on acoustic recordings. Journal of Open Research Software, 5(1):6–??, February 21, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
154/.

Baddour:2017:MCD

[BC17] Natalie Baddour and Ugo Chouinard. Matlab code for the Discrete Hankel Transform. Journal of Open Research Software, 5(1):4–??, January 11, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
82/.

Boettiger:2015:BSB

[BCHR15] Carl Boettiger, Scott Chamberlain, Edmund Hart, and Karthik Ram. Building software, building community: Lessons from the rOpenSci project. Journal of Open Research Software, 3(1):e8–??, November 16, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bu/.

Bisotti:2018:PFF

[BCOP^{+}18] Marc-Antonio Bisotti, David Cortés-Ortuño, Ryan Pepper, Weiwei Wang, Marijan Beg, Thomas Kluyver, and Hans Fangohr. Fidimag — a finite difference atomistic and micromagnetic simulation package. Journal of Open Research Software, 6(1):22–??, September 06, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
223/.

Bethune:2015:PGP

[BG15] Iain Bethune and Yves Gallot. Genefer: Programs for finding large probable generalized Fermat primes. Journal of Open Research Software, 3(1):e10–??, November 19, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ca/.

Brenner:2019:SIW

[BGB^{+}19] Johannes Brenner, Guilio Genova, Giacomo Bertoldi, Georg Niedrist, and Stefano Della

Chiesa. SWCalibrateR: In-

teractive, Web-based calibration of soil moisture sensors. Journal of Open Research Software, 7(1):20–??, June 20, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
254/.

Beckham:2016:PWC

[BHF16] Christopher Beckham, Mark Hall, and Eibe Frank. WekaPyScript:

Classification, regression, and filter schemes for WEKA implemented in Python. Journal of Open Research Software, 4(1):e33–??, August 08, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
108/.

Brown:2018:PPP

[BHS18] Thomas Brown, Jonas Hörsch, and David Schlachtberger. PyPSA: Python for power system analysis. Journal of Open Research Software, 6 (1):4–??, January 16, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
188/.

Blanton:2014:SPS

[BL14] Brian Blanton and Chris Lenhardt. A scientist’s perspective on sustainable scientific software. Journal of Open Research Software, 2 (1):e17–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ba/.

Bond-Lamberty:2019:GRP

[BLDC^{+}19] Ben Bond-Lamberty, Kalyn Dorheim, Ryna Cui, Russell Horowitz, Abigail Snyder, Katherine Calvin, Leyang Feng, Rachel Hoesly, Jill Horing, G. Page Kyle, Robert Link, Pralit Patel, Christopher Roney, Aaron Staniszewski, Sean Turner, Min Chen, Felip Feijoo, Corinne Hartin, Mohamad Hejazi, Gokul Iyer, Sonny Kim, Yaling Liu, Cary Lynch, Haewon McJeon, Steven Smith, Stephanie Waldhoff, Marshall Wise, and Leon Clarke. gcamdata: an R package for preparation, synthesis, and tracking of input data for the GCAM integrated human– Earth systems model. Journal of Open Research Software, 7(1):6–??, March 14, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
232/.

Barnkob:2021:DMT

[BR21] Rune Barnkob and Massimil-

iano Rossi. DefocusTracker: a

modular toolbox for defocusingbased, single-camera, 3D particle tracking. Journal of Open Research Software, 9 (1):22–??, July 23, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
351/.

Broman:2014:FYP

[Bro14] Karl Broman. Fourteen years of R/qtl: Just barely sustainable. Journal of Open Research Software, 2(1):e11–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.

com/articles/10.5334/jors.
at/.

Berrada:2022:RAA

[BS22] Meryem Berrada and Richard A.

Secco. Rho: Application to analyze electrical resistivity. Journal of Open Research Software, 10(1):??, ???? 2022. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
417.

Barbeau:2018:PCO

[BVL18] Xavier Barbeau, Antony T.

Vincent, and Patrick Lagüe. ConfBuster: Open-source tools for macrocycle conformational search and analysis. Journal of Open Research Software, 6(1):1–??, January 08, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
189/.

Bryant:2016:PWW

[BW16] Nathan Bryant and Jeremy Wildfire. Webcharts — a Webbased charting library for custom interactive data visualization. Journal of Open Research Software, 4(1):e29–??, July 19, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Barry:2021:GLE

[BZSH21] Dan Barry, Qijian Zhang, Pheobe Wenyi Sun, and Andrew Hines. Go Listen: an end-to-end online listening test platform. Journal of Open Research Software, 9(1):20–??, July 20, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        361/.

Crusoe:2016:CCC

[CB16a] Michael Crusoe and C. Brown. Channeling community contributions to scientific software: A Sprint experience. Journal of Open Research Software, 4(1):e27–??, July 19, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        96/.

Crusoe:2016:WTA

[CB16b] Michael Crusoe and C. Brown. Walking the talk: Adopting and adapting sustainable scientific software development processes in a small biology lab. Journal of Open Research Software, 4(1):e44–??, November 29, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        35/.

Cheung:2020:FOW

[CC20] Foo Cheung and The CHI Consortium. A figure one Web tool for visualization of experimental designs. Journal of Open Research Software, 8(1):6–??, March 30, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
243/.

Cohen:2014:SDU

[CCH^{+}14] Jeremy Cohen, Chris Cantwell, Neil Chue Hong, David Moxey, Malcolm Illingworth, Andrew Turner, John Darlington, and Spencer Sherwin. Simplifying the development, use and sustainability of HPC software. Journal of Open Research Software, 2(1):e16–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
az/.

Carlsson:2019:TJT

[CE19] Kristoffer Carlsson and Fredrik Ekre. Tensors.jl — tensor computations in Julia. Journal of Open Research Software, 7(1):7–??, March 21, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
182/.

Cheung:2017:WTN

[CFC^{+}17] Foo Cheung, Giovanna Fantoni, Maria Conner, Brian A. Sellers, Yuri Kotliarov, Julián Candia, Katherine Stagliano, and Angélique Biancotto. Web tool for navigating and plotting SomaLogic ADAT files. Journal of Open Research Software, 5(1):20–??, September 08, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
166/.

Crick:2017:RRS

[CHI17] Tom Crick, Benjamin Hall, and Samin Ishtiaq. Reproducibility in research: Systems, infrastructure, culture. Journal of Open Research Software, 5(1):32–??, November 09, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
73/.

Choi:2014:PMQ

[Cho14] Sou-Cheng Choi. MINRES-QLP Pack and reliable reproducible research via supportable scientific software. Journal of Open Research Software, 2(1):e22–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bb/.

Chowdhury:2020:PNC

[CKDG20] A. F. M. Kamal Chowdhury, Jordan Kern, Thanh Duc Dang, and Stefano Galelli. PowNet: a network-constrained unit commitment/economic dispatch model for large-scale power systems analysis. Journal of Open Research Software, 8(1):5–??, March 12, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
302/.

Calvin:2019:GVR

[CLW19] Katherine Calvin, Robert Link, and Marshall Wise. gcamland v1.0 — an R package for modelling land use and land cover change. Journal of Open Research Software, 7(1):31–??, October 22, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
233/.

Cross:2021:SRI

[CNS^{+}21] Tiger Cross, Rasika Navarange, Joon-Ho Son, William Burr, Arjun Singh, Kelvin Zhang, Miruna Rusu, Konstantinos Gkoutzis, Andrew Osborne, and Bart Nieuwenhuis. Simple RGC: ImageJ plugins for counting retinal ganglion cells and determining the transduction efficiency of viral vectors in retinal wholemounts. Journal of Open Research Software, 9(1):15–??, June 09, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Coelho:2013:PMO

[Coe13] Luis Pedro Coelho. Mahotas: Open source software for scriptable computer vision. Journal of Open Research Software, 1(1):e3–??, July 29, 2013. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        ac/.

Coelho:2017:PJS

[Coe17] Luis Pedro Coelho. Jug: Software for parallel reproducible computation in Python. Journal of Open Research Software, 5(1):30–??, October 27, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        161/.

Cresson:2022:STS

[Cre22] Rémi Cresson. SR4RS: a tool for super resolution of remote sensing images. Journal of Open Research Software, 10(1):??, March 02, 2022. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        369/.

Clerx:2019:PIN

[CRL^{+}19] Michael Clerx, Martin Robinson, Ben Lambert, Chon Lok Lei, Sanmitra Ghosh, Gary R. Mirams, and David J. Gavaghan. Probabilistic Inference on Noisy Time Series (PINTS).

Journal of Open Research Software, 7(1):23–??, July 19, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
252/.

Champseix:2021:PPH

[CRL21] Robin Champseix, Laurent Ribiere, and Clément Le Couedic. A Python package for heart rate variability analysis and signal preprocessing. Journal of Open Research Software, 9(1):28–??, October 06, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
305/.

Christ:2023:PJU

[CSR^{+}23] Simon Christ, Daniel Schwabeneder, Christopher Rackauckas, Michael Krabbe Borregaard, and Thomas Breloff. Plots.jl — a user extendable plotting API for the Julia programming language. Journal of Open Research Software, 11(1):??, ???? 2023. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
431.

Chen:2019:PPL

[CZZ19] Zhangqi Chen, Qiaofu Zhang, and Ji-Cheng Zhao. pydiffusion: a Python library for diffusion simulation and data analysis. Journal of Open Research Software, 7(1):13–??, April 23, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        255/.

Dowrick:2020:MSF

[DAF^{+}20] Thomas Dowrick, James Avery, Mayo Faulkner, David Holder, and Kirill Aristovich. EIT-MESHER — segmented FEM mesh generation and refinement. Journal of Open Research Software, 8(1):27–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        321/.

Dowrick:2021:CCT

[DAT^{+}21] Thomas Dowrick, Mian Ahmad, Stephen Thompson, James Hetherington, Jonathan Cooper, and Matt Clark-

                                        son. CMakeCatchTemplate: a

C++ template project. Journal of Open Research Software, 9(1):17–??, July 16, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        319/.

Dawson:2016:PEL

[Daw16a] Andrew Dawson. eofs: A library for EOF analysis of meteorological, oceanographic, and climate data. Journal of Open Research Software, 4(1):e14–??, April 26, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
122/.

Dawson:2016:PWH

[Daw16b] Andrew Dawson. Windspharm:

A high-level library for global wind field computations using spherical harmonics. Journal of Open Research Software, 4(1):e31–??, August 02, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
129/.

Dovlo:2015:TCF

[DB15] Edem Dovlo and Natalie Baddour. Toolbox for the computation of 2D Fourier transforms in polar coordinates via Maple. Journal of Open Research Software, 3(1):e3–??, February 05, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bo/.

Diem:2017:PVP

[DB17] Alexandra K. Diem and Neil W. Bressloff. VaMpy: A Python package to solve 1D blood flow problems. Journal of Open Research Software, 5(1):17–??, June 08, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Dubey:2014:SAM

[DBB^{+}14] Anshu Dubey, Steve Brandt, Richard Brower, Merle Giles, Paul Hovland, Donald Lamb, Frank Löffler, Boyana Norris, Brian O’Shea, Claudio Rebbi, Marc Snir, Rajeev Thakur, and Petros Tzeferacos. Software abstractions and methodologies for HPC simulation codes on future architectures. Journal of Open Research Software, 2(1):e14–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        aw/.

deBuyl:2017:PRS

[dBHD17] Pierre de Buyl, Mu-Jie Huang, and Laurens Deprez. RMPCDMD: Simulations of colloids with coarse-grained hydrodynamics, chemical reactions and external fields. Journal of Open Research Software, 5 (1):3–??, January 11, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        142/.

Dunbar:2021:BPP

[DF21] M. Bekker-Nielsen Dunbar and Thomas J. R. Finnie. bayesint: a Python package for calculating Bayesian credible intervals of ratios of beta distributions. Journal of Open Research Software, 9(1):35–??, December 21, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
283/.

Defrance:2016:PMN

[DLA^{+}16] Josselin Defrance, Caroline Lemaître, Rabih Ajib, Jessica Benedicto, Emilien Mallet, Rémi Pollès, Jean-Pierre Plumey, Martine Mihailovic, Emmanuel Centeno, Cristian Ciracì, David Smith, and Antoine Moreau. Moosh: A numerical Swiss Army knife for the optics of multilayers in Octave/Matlab. Journal of Open Research Software, 4(1):e13–??, April 22, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
100/.

Downs:2015:CRS

[DLR^{+}15] Robert R. Downs, W. Christopher Lenhardt, Erin Robinson, Ethan Davis, and Nicholas Weber. Community recommendations for sustainable scientific software. Journal of Open Research Software, 3(1):e11–??, November 18, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bt/.

Daerr:2016:PPI

[DM16] Adrian Daerr and Adrien

Mogne. Pendent Drop: An

ImageJ plugin to measure the surface tension from an image of a pendent drop. Journal of Open Research Software, 4(1):e3–??, January 28, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        97/.

Damerow:2017:GES

[DPL17] Julia Damerow, B. R. Erick Peirson, and Manfred D. Laubichler. The Giles Ecosystem — storage, text extraction, and OCR of documents. Journal of Open Research Software, 5(1):26–??, September 28, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        164/.

Demski:2016:PEK

[DS16] Andrés Demski and Mariano Llamedo Soria. ecg-kit: a Matlab toolbox for cardiovascular signal processing. Journal of Open Research Software, 4(1):e8–??, April 04, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        86/.

Dubey:2014:ESE

[DV14] Anshu Dubey and Brian Van Straalen. Experiences from software engineering of large scale AMR multiphysics code frameworks. Journal of Open Research Software, 2(1):e7–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
am/.

DiVittorio:2020:MVD

[DVS20] Alan V. Di Vittorio, Chris R.

Vernon, and Shijie Shu. Moirai version 3: a data processing system to generate recent historical land inputs for global modeling applications at various scales. Journal of Open Research Software, 8(1):1–??, January 15, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
266/.

Elnesr:2017:PNN

[EA17] Mohammad N. Elnesr and

A. A. Alazba. NeuroCharter:

A neural networks software to visually discover the effects and contributions between interrelated features. Journal of Open Research Software, 5(1):19–??, September 04, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
135/.

Eerland:2017:PTP

[EBFS17] Willem Eerland, Simon Box, Hans Fangohr, and András Sóbester. Teetool — a probabilistic trajectory analysis tool. Journal of Open Research Software, 5(1):14–??, May 17, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
163/.

Eerland:2017:CRS

[EBS17] Willem Eerland, Simon Box, and András Sóbester. Cambridge rocketry simulator — a stochastic six-degrees-of-freedom rocket flight simulator. Journal of Open Research Software, 5(1):5–??, February 21, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
137/.

Emsley:2019:PTA

[EGB^{+}19] Iain Emsley, Torø Graven, Nicola Bird, Susan Griffiths, Jessica Suess, and Lucy Shaw. Please touch the art: Experiences in developing for the visually impaired. Journal of Open Research Software, 7(1):4–??, March 13, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
231/.

Edhlund:2021:POO

[EML21] Ian Edhlund, Matthew Macauley, and Cindy Lee. PBTK optimizer: an open source application for PBTK model parameter optimization in Python. Journal of Open Research Software, 9(1):4–??, April 14, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
285/.

Edwards:2021:BRP

[ES21] Brandon P. M. Edwards and Adam C. Smith. bbsBayes: an R package for hierarchical Bayesian analysis of North American breeding bird survey data. Journal of Open Research Software, 9(1):19–??, July 20, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
329/.

Farrell:2016:PDA

[Far16] Damien Farrell. DataExplore:

An application for general data analysis in research and education. Journal of Open Research Software, 4(1):e9–??, March 22, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
94/.

Fisher:2020:USM

[FCY20] Paige Fisher, Wentao Cao, and Qingzhao Yu. Using SAS macros for multiple mediation analysis in R. Journal of Open Research Software, 8(1):22–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        277/.

Ferrari:2021:CRP

[Fer21] Diogo Ferrari. CESER: an R package to compute cluster estimated standard errors. Journal of Open Research Software, 9(1):32–??, November 30, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        355/.

Fehr:2021:SRS

[FHRS21] J. Fehr, C. Himpe, S. Rave, and J. Saak. Sustainable research software hand-over. Journal of Open Research Software, 9(1):5–??, April 30, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        307/.

Fachada:2016:PTS

[FLMR16a] Nuno Fachada, Vitor Lopes, Rui Martins, and Agostinho Rosa. PerfAndPubTools — tools for software performance analysis and publishing of results. Journal of Open Research Software, 4(1):e18–??, May 12, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        115/.

Fachada:2016:SUA

[FLMR16b] Nuno Fachada, Vitor Lopes, Rui Martins, and Agostinho Rosa. SimOutUtils — utilities for analyzing time series simulation output. Journal of Open Research Software, 4(1):e38–??, October 21, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
110/.

FiguerasiVentura:2020:PRT

[FLS^{+}20] Jordi Figueras i Ventura, Martin Lainer, Zaira Schauwecker, Jacopo Grazioli, and Urs Germann. Pyrad: a real-time weather radar data processing framework based on Py-ART. Journal of Open Research Software, 8(1):28–??, October 08, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
330/.

Forster:2014:PLW

[FPF14] Matthew Forster, Lyn Paterson, and Robert Forster. LIIS: A web-based system for culture collections and sample annotation. Journal of Open Research Software, 2(1):e2–??, April 08, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
aj/.

Foose:2016:PVF

[FS16] Daniel P. Foose and Ioana E. P. Sizemore. Vespucci: A free, cross-platform tool for spectroscopic data analysis and imaging. Journal of Open Research Software, 4(1):e4–??, January 29, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        91/.

Fulop:2019:VSS

[FS19] Sean A. Fulop and Hannah Scott. Vowel system sandbox: Complex system modelling of language change. Journal of Open Research Software, 7(1):8–??, March 25, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        198/.

Flam-Shepherd:2018:PVP

[FSBK18] Rubens Flam-Shepherd, Dev T. Britto, and Herbert J. Kronzucker. vaCATE: A platform for automating data output from compartmental analysis by tracer efflux. Journal of Open Research Software, 6(1):21–??, August 17, 2018. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        175/.

Fomel:2013:PMO

[FSV^{+}13] Sergey Fomel, Paul Sava, Ioan Vlad, Yang Liu, and Vladimir Bashkardin. Madagascar: open-source software project for multidimensional data analysis and reproducible computational experiments. Journal of Open Research Software, 1(1):e8–??, November 21, 2013. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ag/.

Fenn:2018:QBH

[FW18] Roger Fenn and Ansgar Wenzel. Quandle and biquandle homology calculation in R. Journal of Open Research Software, 6(1):3–??, January 16, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
53/.

Graziotin:2013:WBM

[GA13] Daniel Graziotin and Pekka Abrahamsson. A Web-based modeling tool for the SE-MAT Essence theory of software engineering. Journal of Open Research Software, 1(1):e4–??, September 02, 2013. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ad/.

GarciadelCastilloyLopez:2019:MNL

[Gar19] Jose Luis García del Castillo y López. Machina.NET: a library for programming and real-time control of industrial robots. Journal of Open Research Software, 7(1):27–??, August 13, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
247/.

Grady:2018:PPP

[GDP18] Maxwell Grady, Zhongwei Dai, and Karsten Pohl. PLEASE: The Python Low-energy Electron Analysis SuitE — enabling rapid analysis of LEEM and LEED data. Journal of Open Research Software, 6(1):7–??, February 05, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
191/.

Gentile:2020:CVS

[Gen20] Gabriele Gentile. CURSAT ver. 2.1: a simple, resamplingbased, program to generate pseudoreplicates of data and calculate rarefaction curves. Journal of Open Research Software, 8(1):17–??, August 21, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
260/.

Gillman:2017:PGG

[Gil17] Matthew S. Gillman. GENESIS — the GENEric SImulation system for modelling state transitions. Journal of Open Research Software, 5(1):24–??, September 20, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
179/.

Girard:2014:PCS

[Gir14] Jeffrey Girard. CARMA: Software for continuous affect rating and media annotation. Journal of Open Research Software, 2(1):e5–??, July 03, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ar/.

Glaser-Opitz:2016:PMS

[GOB16] Henrich Glaser-Opitz and Kristína Budajová. MATH: A scientific tool for numerical methods calculation and visualization. Journal of Open Research Software, 4(1):e6–??, February 10, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
55/.

Govan:2016:PED

[Gov16] Paul Govan. eAnalytics: Dynamic Web-based analytics for the energy industry. Journal of Open Research Software, 4(1):e45–??, November 30, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.

                                        com/articles/10.5334/jors.
                                        144/.

Gonzalez-Perez:2020:PPP

[GPKL^{+}20] V. Gonzalez-Perez, P. Keil, Y. Li, A. Zülke, R. Burrel, D. Csala, and H. Hoster. A Python package to preprocess the data produced by novonix high-precision battery-testers. Journal of Open Research Software, 8(1):3–??, March 04, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        281/.

Gaston:2014:CIC

[GPP^{+}14] Derek Gaston, John Peterson, Cody Permann, David Andrs, Andrew Slaughter, and Jason Miller. Continuous integration for concurrent computational framework and application development. Journal of Open Research Software, 2(1):e10–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        as/.

Griffin:2018:PPO

[Gri18] William Griffin. pyObs: Opensource software for computerassisted behavioral observation coding. Journal of Open Research Software, 6 (1):5–??, January 26, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
183/.

Gromping:2020:DMR

[Grö20] Ulrike Grömping. DoE.MIParray:

an R package for algorithmic creation of orthogonal arrays. Journal of Open Research Software, 8(1):24–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
286/.

Gebbie-Rayet:2016:PLL

[GRSLL16] James Gebbie-Rayet, Gareth Shannon, Hannes H. Loeffler, and Charles A. Laughton. Longbow: A lightweight remote job submission tool. Journal of Open Research Software, 4(1):e1–??, January 27, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
95/.

Groen:2013:PML

[GRZ13] Derek Groen, Steven Rieder, and Simon Portegies Zwart. MPWide: a light-weight library for efficient message passing over wide area networks. Journal of Open Research Software, 1(1):e9–??, December 20, 2013. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ah/.

Hansen:2016:PPP

[Han16]  Mads Hansen.  PyVDT: A

PsychoPy-based visual sequence detection task. Journal of Open Research Software, 4(1):e22–??, June 10, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        117/.

Han:2021:BIB

[Han21] Hyemin Han. BayesFactorFMRI: Implementing Bayesian secondlevel fMRI analysis with multiple comparison correction and Bayesian meta-analysis of fMRI images with multiprocessing. Journal of Open Research Software, 9 (1):1–??, February 02, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        328/.

Hansen:2022:NFA

[Han22] Johannes N. Hansen. nd — a framework for the analysis of n-dimensional Earth Observation Data. Journal of Open Research Software, 10(1):??, March 11, 2022. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        377/.

Hazra:2014:PIE

[Haz14]  Abhisek Hazra.  Image

Enhancer: A graphic editor to apply numerous effects in digital image. Journal of Open Research Software, 2(1):e31–??, December 11, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bm/.

Heckmann:2017:PGS

[HB17] Mark Heckmann and Lukas

Burk. Gridsampler — a

simulation tool to determine the required sample size for repertory grid studies. Journal of Open Research Software, 5(1):2–??, January 10, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
150/.

Helmus:2016:PAR

[HC16] Jonathan Helmus and Scott Collis. The Python ARM Radar Toolkit (Py-ART), a library for working with weather radar data in the Python programming language. Journal of Open Research Software, 4(1):e25–??, July 18, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
119/.

Herring:2021:PMD

[HFM^{+}21] Angela Herring, Charles Ferenbaugh, Christopher Malone, Daniel Shevitz, Evgeny Kikinzon, Gary Dilts, Hoby Rakotoarivelo, Jan Velechovsky, Konstantin Lipnikov, Navamita Ray, and Rao Garimella. Portage: a modular data remap library for multiphysics applications on advanced architectures. Journal of Open Research Software, 9(1):26–??, September 28, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
338/.

Hagdorn:2023:TCS

[HG23] Magnus Hagdorn and Noel Gourmelen. Taskfarm: a client/server framework for supporting massive embarrassingly parallel workloads. Journal of Open Research Software, 11(1):??, ???? 2023. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
393.

Hadjimichael:2020:RPL

[HGHR20] Antonia Hadjimichael, David Gold, David Hadka, and Patrick Reed. Rhodium: Python library for manyobjective robust decision making and exploratory modeling. Journal of Open Research Software, 8(1):12–??, June 09, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.

com/articles/10.5334/jors.
293/.

Hilpert:2021:OTI

[HGS21] Simon Hilpert, Stephan Günther, and Martin Söthe. oemof.tabular — introducing data packages for reproducible workflows in energy system modeling. Journal of Open Research Software, 9(1):6–??, May 05, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
320/.

Haeri:2015:IPA

[HH15] Morteza Haeri and Mohammad Haeri. ImageJ plugin for analysis of porous scaffolds used in tissue engineering. Journal of Open Research Software, 3(1):e1–??, January 28, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bn/.

Hoyer:2017:PXL

[HH17] Stephan Hoyer and Joe Hamman. xarray: N-D labeled arrays and datasets in Python. Journal of Open Research Software, 5(1):10–??, April 05, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Hussan:2020:CSS

[HH20] Jagir R. Hussan and Peter J. Hunter. Comfort Simulator: a software tool to model thermoregulation and perception of comfort. Journal of Open Research Software, 8(1):16–??, July 20, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
288/.

Hoffmann:2021:PPC

[HKGvS21] Christoph G. Hoffmann, George N. Kiladis, Maria Gehne, and Christian von Savigny. A Python package to calculate the OLR-based index of the Madden–Julian-Oscillation (OMI) in climate science and weather forecasting. Journal of Open Research Software, 9(1):9–??, May 14, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
331/.

Hughes:2015:PSS

[HLR15] Adam Hughes, Zhaowen Liu,

and M. Reeves. Scikit-spectra:

Explorative spectroscopy in Python. Journal of Open Research Software, 3(1):e6–??, June 05, 2015. CODEN ???? ISSN 2049-9647. URL https:/

/openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bs/.

Hussain:2023:PPP

[HMB23] Md. Manjurul Hussain, Ishtiak Mahmud, and Sheikh Hefzul

Bari.  pyHomogeneity: a

Python package for homogeneity test of time series data. Journal of Open Research Software, 11(1):??, ???? 2023. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
427.

Henderson:2017:PPM

[HR17] Ryan Henderson and Rasmus Rothe. Picasso: A modular framework for visualizing the learning process of neural network image classifiers. Journal of Open Research Software, 5(1):22–??, September 11, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
178/.

Heck:2018:DOS

[HS18] Nina Heck and Moritz Schubotz.

DiViDu — an open source solution for dual task experiments with integrated divided visual field paradigm. Journal of Open Research Software, 6(1):14–??, April 06, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Hughes:2014:PPI

[Hug14]  Adam Hughes.  pyparty:

Intuitive particle processing in Python. Journal of Open Research Software, 2 (1):e26–??, September 23, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        bh/.

Harvey:2019:TPS

[HW19] Christine Harvey and R. S.

                                        Weigel. Transplant2Mongo:

Python scripts that insert organ procurement and transplantation network (OPTN) data in MongoDB. Journal of Open Research Software, 7(1):5–??, March 14, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        229/.

Huang:2016:POT

[HWX^{+}16] Shuai Huang, Yaqiang Wang, Yuanli Xie, Peng Zhao, and Johannes Lüers. OutlierFlag: A tool for scientific data quality control by outlier data flagging. Journal of Open Research Software, 4(1):e20–??, May 31, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        90/.

Hewitt:2016:MWA

[HYG16] Brett Hewitt, Moi Hoon Yap, and Robyn Grant. Manual Whisker Annotator (MWA): A modular open-source tool. Journal of Open Research Software, 4(1):e16–??, April 28, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
93/.

Illing:2016:CPM

[IKKC16] Sebastian Illing, Christopher Kadow, Oliver Kunst, and Ulrich Cubasch. Correction: MurCSS: A tool for standardized evaluation of decadal hindcast systems. Journal of Open Research Software, 4(1):e24–??, June 24, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

136/. See [IKOC14].

Illing:2014:PMT

[IKOC14] Sebastian Illing, Christopher Kadow, Kunst Oliver, and Ulrich Cubasch. MurCSS: A tool for standardized evaluation of decadal hindcast systems. Journal of Open Research Software, 2(1):e24–??, September 04, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

bf/. See correction [IKKC16].

Jacobs:2014:PPP

[JAGP14] Christian Jacobs, Alexandros Avdis, Gerard Gorman, and Matthew Piggott. PyRDM: A Python-based library for automating the management and online publication of scientific software and data. Journal of Open Research Software, 2(1):e28–??, October 03, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        bj/.

Jayasena:2021:RTL

[Jay21]  Aruna Jayasena.  Register

transfer level disparity generator with stereo vision. Journal of Open Research Software, 9(1):18–??, July 19, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        339/.

Jackson:2020:PPD

[JCL^{+}20] Robert Jackson, Scott Collis, Timothy Lang, Corey Potvin, and Todd Munson. PyDDA: a Pythonic direct data assimilation framework for wind retrievals. Journal of Open Research Software, 8(1):20–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        264/.

Jelinek:2021:RTL

[JLZ21] Ales Jelinek, Adam Ligocki, and Ludek Zalud. Robotic template library. Journal of Open Research Software, 9(1):25–??, September 20, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
353/.

Jacobsen:2017:GSJ

[JNR17] Robert Dahl Jacobsen, Morten Nielsen, and Morten Grud Rasmussen. Generalized sampling in Julia. Journal of Open Research Software, 5(1):12–??, April 20, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
157/.

Jones:2018:PMM

[Jon18] Pete Richard Jones. Myex: A MATLAB interface for the Tobii Eyex eye-tracker. Journal of Open Research Software, 6(1):16–??, April 20, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
196/.

Jucker:2014:SVA

[Juc14]  Martin Jucker.  Scientific

visualisation of atmospheric data with ParaView. Journal of Open Research Software, 2(1):e4–??, June 03, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.

com/articles/10.5334/jors.
al/.

Katz:2014:TCM

[Kat14] Daniel Katz. Transitive credit as a means to address social and technological concerns stemming from citation and attribution of digital products. Journal of Open Research Software, 2(1):e20–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
be/.

Knight:2016:OFR

[KCH^{+}16] Vincent Knight, Owen Campbell, Marc Harper, Karol Langner, James Campbell, Thomas Campbell, Alex Carney, Martin Chorley, Cameron Davidson-Pilon, Kristian Glass, Nikoleta Glynatsi, Tomás Ehrlich, Martin Jones, Georgios Koutsovoulos, Holly Tibble, Jochen Müller, Geraint Palmer, Piotr Petunov, Paul Slavin, Timothy Standen, Luis Visintini, and Karl Molden. An open framework for the reproducible study of the Iterated Prisoner’s Dilemma. Journal of Open Research Software, 4(1):e35–??, August 31, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
125/.

Katz:2014:SFW

[KCL^{+}14] Daniel Katz, Sou-Cheng Choi, Hilmar Lapp, Ketan Maheshwari, Frank Löffler, Matthew Turk, Marcus Hanwell, Nancy Wilkins-Diehr, James Hetherington, James Howison, Shel Swenson, Gabrielle Allen, Anne Elster, Bruce Berriman, and Colin Venters. Summary of the First Workshop on Sustainable Software for Science: Practice and Experiences (WSSSPE1). Journal of Open Research Software, 2(1):e6–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
an/.

Katz:2016:RTW

[KCN^{+}16] Daniel Katz, Sou-Cheng Choi, Kyle Niemeyer, James Hetherington, Frank Löffler, Dan Gunter, Ray Idaszak, Steven Brandt, Mark Miller, Sandra Gessing, Nick Jones, Nic Weber, Suresh Marru, Gabrielle Allen, Birgit Penzenstadler, Colin Venters, Ethan Davis, Lorraine Hwang, Ilian Todorov, Abani Patra, and Miguel de Val-Borro. Report on the Third Workshop on Sustainable Software for Science: Practice and Experiences (WSSSPE3). Journal of Open Research Software, 4(1):e37–??, October 21, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Katz:2016:RSW

[KCWD^{+}16] Daniel S. Katz, Sou-Cheng T. Choi, Nancy Wilkins-Diehr, Neil Chue Hong, Colin C. Venters, James Howison, Frank Seinstra, Matthew Jones, Karen A. Cranston, Thomas L. Clune, Miguel de Val-Borro, and Richard Littauer. Report on the Second Workshop on Sustainable Software for Science: Practice and Experiences (WSSSPE2). Journal of Open Research Software, 4(1):e7–??, February 22, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        85/.

Katz:2019:SSR

[KDH^{+}19] Daniel S. Katz, Stephan Druskat, Robert Haines, Caroline Jay, and Alexander Struck. The state of sustainable research software: Learning from the workshop on sustainable software for science: Practice and experiences (WSSSPE5.1). Journal of Open Research Software, 7(1):11–??, April 02, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        242/.

Kelter:2020:BRP

[Kel20] Riko Kelter. bayest: an R package for effect-size targeted Bayesian two-sample t-tests. Journal of Open Research Software, 8(1):14–??, June 15, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
290/.

Kaiser:2020:JPP

[KFV20] K. E. Kaiser, A. N. Flores, and C. R. Vernon. Janus: a Python package for agentbased modeling of land use and land cover change. Journal of Open Research Software, 8(1):15–??, June 25, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
306/.

Korosov:2016:PNS

[KHD^{+}16] Anton Korosov, Morten Hansen, Knut-Frode Dagestad, Asuka Yamakawa, Aleksander Vines, and Maik Riechert. Nansat: a scientist-orientated Python package for geospatial data processing. Journal of Open Research Software, 4 (1):e39–??, October 24, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
120/.

Kibele:2016:BPS

[Kib16] Jared Kibele. Benthic photo survey: Software for geotagging, depth-tagging, and classifying photos from survey data and producing shapefiles for habitat mapping in GIS. Journal of Open Research Software, 4(1):e10–??, March 23, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
104/.

Kadow:2021:IFF

[KILE^{+}21] Christopher Kadow, Sebastian Illing, Etor E. Lucio-Eceiza, Martin Bergemann, Mahesh Ramadoss, Philipp S. Sommer, Oliver Kunst, Thomas Schartner, Klaus Pankatz, Jens Grieger, Mareike Schuster, Andy Richling, Hannes Thiemann, Ingo Kirchner, Henning W Rust, Thomas Ludwig, Ulrich Cubasch, and Uwe Ulbrich. Introduction to Freva — a free evaluation system framework for Earth system modeling. Journal of Open Research Software, 9(1):13–??, June 02, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
253/.

Kimbrough:2017:PAC

[Kim17] Gray Kimbrough. ATUS-Commuting: Stata code to calculate commuting time in the American Time Use Survey. Journal of Open Research Software, 5(1):25–??, September 25, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
162/.

Kotsalos:2021:PNS

[KLC21] Christos Kotsalos, Jonas Latt, and Bastien Chopard. Palabos-npFEM: Software for the simulation of cellular blood flow (digital blood). Journal of Open Research Software, 9 (1):16–??, June 24, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
343/.

Katz:2018:FWS

[KNG^{+}18] Daniel S. Katz, Kyle E.

Niemeyer, Sandra Gesing, Lorraine Hwang, Wolfgang Bangerth, Simon Hettrick, Ray Idaszak, Jean Salac, Neil Chue Hong, Santiago Núñez-Corrales, Alice Allen, R. Stuart Geiger, Jonah Miller, Emily Chen, Anshu Dubey, and Patricia Lago. Fourth Workshop on Sustainable Software for Science: Practice and Experiences (WSSSPE4). Journal of Open Research Software, 6(1):10–??, February 15, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
184/.

Katz:2014:FDR

[KP14] Daniel Katz and David Proctor. A framework for discussing e-research infrastructure sustainability. Journal of Open Research Software, 2(1):e13–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        av/.

Krakauer:2016:PSP

[Kra16]  Nir Krakauer.  SeFo: A

package for generating probabilistic forecasts from NMME predictive ensembles. Journal of Open Research Software, 4(1):e19–??, May 20, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        112/.

Katz:2015:TCP

[KS15] Daniel S. Katz and Arfon M. Smith. Transitive credit and JSON-LD. Journal of Open Research Software, 3(1):e7–??, November 05, 2015. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        by/.

Khan:2020:MTH

[KWV^{+}20] Zarrar Khan, Thomas Wild, Chris Vernon, Andy Miller, Mohamad Hejazi, Leon Clarke, Fernando Miralles-Wilhelm, Raul Munoz Castillo, Fekadu Moreda, Julia Lacal Bereslawski, Micaela Suriano, and Jose Casado. Metis — a tool to harmonize and analyze multisectoral data and linkages at variable spatial scales. Journal of Open Research Software, 8(1):10–??, March 30, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
292/.

Lenhardt:2014:DML

[LAB^{+}14] W Lenhardt, Stanley Ahalt, Brian Blanton, Laura Christopherson, and Ray Idaszak. Data management lifecycle and software lifecycle management in the context of conducting science. Journal of Open Research Software, 2(1):e15–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ax/.

Lawton:2017:DSD

[Law17]  Tom Lawton.  DSCImage-

Calc — software for determining similarity coefficients for the analysis of image segmentations. Journal of Open Research Software, 5(1):28–??, October 18, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
16/.

Loffler:2014:PCI

[LBAS14] Frank Löffler, Steven Brandt, Gabrielle Allen, and Erik Schnetter. Cactus: Issues for sustainable simulation software. Journal of Open Research Software, 2 (1):e12–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        au/.

Lee:2019:CII

[LD19] Edward D. Lee and Bryan C. Daniels. Convenient Interface to Inverse Ising (ConIII): a Python 3 package for solving Ising-type maximum entropy models. Journal of Open Research Software, 7 (1):3–??, March 04, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        217/.

Lima:2022:UUN

[LFD^{+}22] Brena Lima, Nuno Fachada, Rui Dinis, Daniel B. Costa, and Marko Beko. Uavnoma: a UAV-NOMA network model under non-ideal conditions. Journal of Open Research Software, 10(1):??, October 17, 2022. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        397/.

Liu:2017:POO

[Liu17] Tianshu Liu. OpenOpticalFlow: An open source program for extraction of velocity fields from flow visualization images. Journal of Open Research Software, 5(1):29–??, October 20, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
168/.

Lamprecht:2022:WDW

[LMOB^{+}22] Anna-Lena Lamprecht, Carlos Martinez-Ortiz, Michelle Barker, Sadie L. Bartholomew, Justin Barton, Neil Chue Hong, Jeremy Cohen, Stephan Druskat, Jeremy Forest, Jean-Noël Grad, Daniel S. Katz, Robin Richardson, Robert Rosca, Douwe Schulte, Alexander Struck, and Marion Weinzierl. What do we (not) know about research software engineering? Journal of Open Research Software, 10(1):??, ???? 2022. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
384.

Lougheed:2016:PMO

[LO16] Bryan Lougheed and Stephen Obrochta. MatCal: Open source Bayesian ^{14}C age calibration in Matlab. Journal of Open Research Software, 4(1):e42–??, November 08, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
130/.

Leightley:2018:PIO

[LPG^{+}18] Daniel Leightley, Jo-Anne Puddephatt, Laura Goodwin, Roberto Rona, and Nicola T.

Fear. InDEx: Open source iOS and Android software for self-reporting and monitoring of alcohol consumption. Journal of Open Research Software, 6(1):13–??, March 23, 2018. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        207/.

Liu:2021:OOS

[LS21] Tianshu Liu and David M.

                                        Salazar. OpenOpticalFlow PIV:

an open source program integrating optical flow method with cross- correlation method for particle image velocimetry. Journal of Open Research Software, 9(1):3–??, February 05, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        326/.

Lukacs:2021:SCS

[Luk21] Gáspár Lukács. SmaRT-CIT: Smartphone app for the response time concealed information test with ready-to-use Android deployment. Journal of Open Research Software, 9(1):21–??, July 21, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        341/.

Luyten:2013:PET

[LV13] Laura Luyten and Frederik Van Cappellen. ExpTimer: timer software to facilitate complex, multi-step procedures. Journal of Open Research Software, 1(1):e2–??, April 14, 2013. CODEN ???? ISSN 2049-9647. URL https:/

/openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ab/.

Li:2017:PXG

[LVH^{+}17] Xinya Li, Chris R. Vernon, Mohamad I. Hejazi, Robert P. Link, Leyang Feng, Yaling Liu, and Lynn T. Rauchenstein. Xanthos — a global hydrologic model. Journal of Open Research Software, 5(1):21–??, September 11, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
181/.

Li:2018:PTP

[LVH^{+}18] Xinya Li, Chris R. Vernon, Mohamad I. Hejazi, Robert P. Link, Zhongwei Huang, Lu Liu, and Leyang Feng. Tethys

— a Python package for spatial and temporal downscaling of global water withdrawals. Journal of Open Research Software, 6(1):9–??, February 09, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
197/.

Lansley:2016:PCI

[LVSF16] Alastair Lansley, Peter Vamplew, Philip Smith, and Cameron Foale. Caliko: An inverse kinematics software library implementation of the FABRIK algorithm. Journal of Open Research Software, 4(1):e36–??, September 09, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        116/.

Lavielle:2020:ECU

[LW20] Marc Lavielle and Philip D. Waggoner. Exploring and comparing unsupervised clustering algorithms. Journal of Open Research Software, 8(1):21–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        269/.

Marciuska:2013:PUE

[MA13] Sarunas Marciuska and Pekka

                                        Abrahamsson. Feature Usage

Explorer: Usage monitoring and visualization tool in HTML5 based applications. Journal of Open Research Software, 1(1):e7–??, October 18, 2013. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        af/.

Markussen:2016:PPC

[Mar16] Thor Markussen. Parchar — characterization of suspended particles through image processing in Matlab. Journal of Open Research Software, 4(1):e26–??, July 19, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
114/.

Matter:2018:PRH

[Mat18]  Ulrich Matter.  RWebData:

A high-level interface to the programmable Web. Journal of Open Research Software, 6(1):11–??, February 21, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
201/.

Mohanan:2019:FCA

[MBA19] Ashwin Vishnu Mohanan, Cyrille Bonamy, and Pierre Augier. FluidFFT: Common API (C++ and Python) for Fast Fourier Transform HPC libraries. Journal of Open Research Software, 7 (1):10–??, April 01, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
238/.

Mohanan:2019:FMO

[MBLA19] Ashwin Vishnu Mohanan, Cyrille Bonamy, Miguel Calpe Linares, and Pierre Augier. FluidSim: Modular, objectoriented Python package for high-performance CFD simulations. Journal of Open Research Software, 7(1):14–??, April 26, 2019. CODEN ???? ISSN 2049-9647. URL https:/

                                        /openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        239/.

Mueller:2014:CTS

[ME14] Shane Mueller and Alena Esposito. Computerized testing software for assessing interference suppression in children and adults: The Bivalent Shape Task (BST). Journal of Open Research Software, 2(1):e3–??, May 23, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        ak/.

Meurs:2016:FRL

[Meu16] Joris Meurs. Flow rates in liquid chromatography, gas chromatography and supercritical fluid chromatography: A tool for optimization. Journal of Open Research Software, 4(1):e32–??, August 02, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        126/.

Moller:2016:MTI

[MGMP16] Birgit Möller, Markus Glaß, Danny Misiak, and Stefan Posch. MiToBo — a toolbox for image processing and analysis. Journal of Open Research Software, 4(1):e17–??, April 28, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
103/.

Michelson:2018:BAW

[MHE^{+}18] Daniel Michelson, Anders Henja, Sander Ernes, Günther Haase, Jarmo Koistinen, Katarzyna O’sródka, Tuomas Peltonen, Maciej Szewczykowski, and Jan Szturc. BALTRAD advanced weather radar networking. Journal of Open Research Software, 6(1):12–??, March 19, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
193/.

Meller:2016:PDM

[ML16] Yosef Meller and Alex Liberzon. Particle data management software for 3D particle tracking velocimetry and related applications — the Flowtracks package. Journal of Open Research Software, 4(1):e23–??, June 16, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
101/.

Molina:2021:HOS

[MOM21] Tulio Molina, Juan Ortega, and Juan Muñoz. HELMpy, open source package of power flow solvers, including the holomorphic embedding load flow method (HELM), developed on Python 3. Journal of Open Research Software, 9(1):23–??, August 18, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        310/.

Moore:2021:AWB

[Moo21] Adrian Moore. autoAPI — a Web-based tool for specification of an API endpoint to return JSON data from an XML source. Journal of Open Research Software, 9(1):24–??, August 20, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        335/.

Mirouze:2021:SSM

[MR21] Isabelle Mirouze and Sophie Ricci. Smurf: System for modelling with uncertainty reduction, and forecasting. Journal of Open Research Software, 9(1):2–??, February 05, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        312/.

Moore:2014:EGD

[MRX14] Reagan Moore, Arcot Rajasekar, and Hao Xu. Extensible generic data management software. Journal of Open Research Software, 2(1):e9–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ap/.

Manning:2014:PCS

[MS14] Saundra Manning and Lior Shamir. CHLOE: A software tool for automatic novelty detection in microscopy image datasets. Journal of Open Research Software, 2(1):e25–??, September 22, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bg/.

Muthusi:2021:SGS

[MYM21] Jacques Muthusi, Peter W.

Young, and Samuel Mwalili. %svy freqs: a generic SAS macro for creating publicationquality three-way cross-tabulations. Journal of Open Research Software, 9(1):30–??, October 22, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
318/.

Nedelec:2017:PPV

[Ned17] François Nedelec. preconfig:

A versatile configuration file generator for varying parameters. Journal of Open Research Software, 5(1):9–??, April 05, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Nguyen:2022:MCT

[NHL^{+}22] Thi-Ngoc-Thu Nguyen, Andreas Höfter, Kevin Leonardic, Stefanie Rosenhain, Fabian Kiessling, Wannida Sae-Tang, Uwe Naumann, and Felix Gremse. A micro-computed tomography database and reference implementation for parallel computations of trabecular thickness and spacing. Journal of Open Research Software, 10(1):??, March 16, 2022. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
360/.

Nikolic:2013:DPT

[NL13] Konstantin Nikolic and Joaquim Loizu. Drosophila phototransduction simulator. Journal of Open Research Software, 1(1):e1–??, April 14, 2013. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/503b9b1a69665/

.

Nygaard:2018:PBT

[NMA^{+}18] Henrik Nygård, Ciarán Murray, Jesper H. Andersen, Georg Martin, Kaire Torn, and Samuli Korpinen. BEAT 3.0

— a tool for integrated biodiversity assessments. Journal of Open Research Software, 6(1):19–??, August 02, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.

com/articles/10.5334/jors.
226/.

Nandakumar:2020:AOC

[NS20] Hari Nandakumar and Shailesh Srivastava. ABC-OCT — a crossplatform implementation of real-time Fourier-domain optical coherence tomography. Journal of Open Research Software, 8(1):7–??, March 30, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
272/.

Osborne:2014:HPO

[OHM14] Nicola Osborne, George Hamilton, and Stuart Macdonald. Historical Post Office Directory Parser (POD parser) software from the AddressingHistory Project. Journal of Open Research Software, 2(1):e23–??, July 21, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
aq/.

Onogi:2016:PVV

[OI16] Akio Onogi and Hiroyoshi Iwata. VIGoR: Variational Bayesian inference for genomewide regression. Journal of Open Research Software, 4 (1):e11–??, April 04, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Oles:2020:BTD

[OK20] Vladyslav Oles and Anton Kukushkin. BoolSi: a tool for distributed simulations and analysis of Boolean networks. Journal of Open Research Software, 8(1):26–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        308/.

OMara:2017:PII

[OKVK17] Aidan O’Mara, Anna E. King, James C. Vickers, and Matthew T. K. Kirkcaldie. ImageSURF: An ImageJ plugin for batch pixelbased image segmentation using random forests. Journal of Open Research Software, 5(1):31–??, November 06, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        172/.

Otis:2017:PPC

[OL17] Richard Otis and Zi-Kui Liu. pycalphad: CALPHADbased computational thermodynamics in Python. Journal of Open Research Software, 5(1):1–??, January 09, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        140/.

Oxvig:2014:PMP

[OPA^{+}14] Christian Oxvig, Patrick Pedersen, Thomas Arildsen, Jan Østergaard, and Torben Larsen. Magni: A Python package for compressive sampling and reconstruction of atomic force microscopy images. Journal of Open Research Software, 2(1):e29–??, October 07, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bk/.

Pastell:2016:CCP

[Pas16] Matti Pastell. CowLog — cross-platform application for coding behaviours from video. Journal of Open Research Software, 4(1):e15–??, April 27, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
113/.

Petrov:2022:OSM

[PAT22] Nikolay Petrov, Vasil Atanasov, and Trevor Thompson. Opensource MUltiple Tests Corrections and FOrmatted Tables Software (MUFOS). Journal of Open Research Software, 10(1):??, March 17, 2022. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Pavesi:2016:ETD

[PBF^{+}16] Fabio Pavesi, Anna Basoni, Cristiano Fugazza, Stefano Menegon, Alessandro Oggioni, Monica Pepe, Paolo Tagliolato, and Paola Carrara. EDI — a template-driven metadata Editor for research data. Journal of Open Research Software, 4(1):e40–??, October 25, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
106/.

Petticrew:2018:SMC

[PDN18] Jonathan Petticrew, Simon Dimler, and Jo Shien Ng. Simple Monte Carlo simulator for modelling linear mode and Geiger mode avalanche photodiodes in C++. Journal of Open Research Software, 6(1):17–??, May 18, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
212/.

Ptaszynski:2017:PMA

[PDR^{+}17] Michał Ptaszyński, Paweł Dybala, Rafał Rzepka, Kenji Araki, and Fumito Masui. ML-Ask: Open source affect analysis software for textual input in Japanese. Journal of Open Research Software, 5(1):16–??, June 07, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.

com/articles/10.5334/jors.
149/.

Plieger:2021:ASI

[PdV21] Maarten Plieger and Ernst

de Vreede. Adaguc-Server:

Interactive access to heterogeneous meteorological and climatological datasets using open standards. Journal of Open Research Software, 9(1):33–??, December 09, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
382/.

Peters:2014:PAE

[Pet14] Nial Peters. AvoPlot: An extensible scientific plotting tool based on matplotlib. Journal of Open Research Software, 2(1):e1–??, February 10, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ai/.

Piasini:2021:EPP

[PFLG21] Eugenio Piasini, Alexandre L. S. Filipowicz, Jonathan Levine, and Joshua I. Gold. Embo: a Python package for empirical data analysis using the information bottleneck. Journal of Open Research Software, 9(1):10–??, May 31, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Paine:2020:EFU

[PGR20] Drew Paine, Devarshi Ghoshal, and Lavanya Ramakrishnan. Experiences with a flexible user research process to build data change tools. Journal of Open Research Software, 8(1):18–??, September 01, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        284/.

Petrowski:2019:RFP

[PKB19] Paul F. Petrowski, Elizabeth G. King, and Timothy M. Beissinger. An R framework for the partitioning of linkage disequilibrium between and within populations. Journal of Open Research Software, 7(1):15–??, April 30, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        250/.

Pluta:2020:EOP

[PL20] Adam Pluta and Ontje Lünsdorf.

                                        esy-osmfilter — a Python

library to efficiently extract OpenStreetMap data. Journal of Open Research Software, 8(1):19–??, September 01, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        317/.

Posch:2017:PAA

[PM17] Stefan Posch and Birgit Moeller. Alida — advanced library for integrated development of data analysis applications. Journal of Open Research Software, 5(1):7–??, March 23, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
124/.

Pierce:2015:PIP

[PMM15] Marlon E. Pierce, Suresh Marru, and Chris Mattmann. Patching it up, pulling it forward. Journal of Open Research Software, 3(1):e12– ??, November 19, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bz/.

Quade:2019:GSR

[QGA19] Markus Quade, Julien Gout, and Markus Abel. Glyph: Symbolic regression tools. Journal of Open Research Software, 7(1):19–??, June 17, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
192/.

Rabe:2020:SMG

[Rab20] Martin Rabe. Spectram: a MATLAB(R) and GNU Octave toolbox for transition model guided deconvolution of dynamic spectroscopic data. Journal of Open Research Software, 8(1):13–??, June 09, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
323/.

Rass:2020:HPP

[Raß20] Alexander Raß. High Precision Particle Swarm Optimization Algorithm (HiPPSO). Journal of Open Research Software, 8(1):4–??, March 09, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
282/.

Radtke:2019:VWB

[RBB^{+}19] Hagen Radtke, Florian Börgel, Sandra-Esther Brunnabend, Anja Eggert, Madline Kniebusch, H. E. Markus Meier, Daniel Neumann, Thomas Neumann, and Manja Placke. Validator — a Web-based interactive tool for validation of ocean models at oceanographic stations. Journal of Open Research Software, 7(1):18–??, June 12, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
259/.

Rosjat:2022:DDS

[RD22] Nils Rosjat and Silvia Daun.

DST (Dynamic Synchronization

Toolbox): a MATLAB implementation of the dynamic phase-locking pipeline from stimulus transformation into motor action: Dynamic graph analysis reveals a posterior-to- anterior shift in brain network communication of older subjects. Journal of Open Research Software, 10(1):??, August 01, 2022. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
394/.

Ravasio:2021:OOO

[RDB21] Claudio S. Ravasio, Lyndon Da Cruz, and Christos Bergeles. oflibnumpy

and oflibpytorch: Opti-

cal flow handling and manipulation in Python. Journal of Open Research Software, 9(1):31–??, November 26, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
380/.

Ramalhinho:2023:FSP

[RDBC23] João Ramalhinho, Thomas Dowrick, Ester Bonmati, and Matthew J. Clarkson. Fan-Slicer: a Pycuda package for fast reslicing of ultrasound shaped planes. Journal of Open Research Software, 11(1):??, ???? 2023. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.

                                        com/articles/10.5334/jors.
                                        422.

Riefstahl:2021:FMP

[RG21] Florian Riefstahl and Felix Gross. FastGAPP — a MAT-LAB based program supports Earth scientists interpreting geochemical, petrological and sedimentological data. Journal of Open Research Software, 9(1):11–??, May 31, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        333/.

Rajananda:2018:RDK

[RLO18] Sivananda Rajananda, Hakwan Lau, and Brian Odegaard. A random-dot kinematogram for Web-based vision research. Journal of Open Research Software, 6(1):6–??, January 27, 2018. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        194/.

Rinn:2016:LAR

[RLWP16] Philip Rinn, Pedro Lind, Matthias Wächter, and Joachim Peinke. The Langevin approach: An R package for modeling Markov processes. Journal of Open Research Software, 4(1):e34–??, August 23, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        123/.

Roberge:2017:PHW

[RML17] Martin C. Roberge, Michael P.

McGuire, and Jie Lian. HydroCloud: A Web application for exploring stream gage data. Journal of Open Research Software, 5 (1):18–??, August 23, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
173/.

Rackauckas:2017:PDJ

[RN17] Christopher Rackauckas and

Qing Nie. DifferentialEquations.jl

— a performant and featurerich ecosystem for solving differential equations in Julia. Journal of Open Research Software, 5(1):15–??, May 25, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
151/.

Raguin:2023:ECT

[RSR23] Adélaïde Raguin, Ian Stansfield, and Maria Carmen Ro-

mano. ExpressInHost: a

codon tuning tool for the expression of recombinant proteins in host microorganisms. Journal of Open Research Software, 11(1):??, ???? 2023. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Richardson:2020:ELV

[RWE^{+}20] Robin A. Richardson, David W. Wright, Wouter Edeling, Vytautas Jancauskas, Jalal Lakhlili, and Peter V. Coveney. EasyVVUQ: a library for verification, validation and uncertainty quantification in high performance computing. Journal of Open Research Software, 8(1):11–??, April 29, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        303/.

Shaad:2020:OST

[SA20] Kashif Shaad and Howard Alt. An open source toolbox for integrating freshwater social-ecological indicators in basin management. Journal of Open Research Software, 8(1):9–??, March 30, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        291/.

Sarra:2017:MRB

[Sar17] Scott A. Sarra. The Matlab Radial Basis Function Toolbox. Journal of Open Research Software, 5(1):8–??, March 27, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        131/.

Schoeffler:2018:WCF

[SBS^{+}18] Michael Schoeffler, Sarah Bartoschek, Fabian-Robert Stöter, Marlene Roess, Susanne Westphal, Bernd Edler, and Jürgen Herre. webMUSHRA — a comprehensive framework for Webbased listening tests. Journal of Open Research Software, 6(1):8–??, February 05, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
187/.

Spencer:2015:OSD

[SBV^{+}15] J. S. Spencer, N. S. Blunt, W. A. Vigor, Fionn D. Malone, W. M. C. Foulkes, James J. Shepherd, and A. J. W. Thom. Open-source development experiences in scientific software: The HANDE Quantum Monte Carlo Project. Journal of Open Research Software, 3(1):e9–??, November 17, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bw/.

Schimpf:2020:CIC

[SC20] Corey Schimpf and Brian Castellani. COMPLEX-IT: a case-based modelling and scenario simulation platform for social inquiry. Journal of Open Research Software, 8(1):25–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        298/.

Schwabe:2017:PBR

[Sch17] Inga Schwabe. BayesTwin: An R package for Bayesian inference of item-level twin data. Journal of Open Research Software, 5(1):33–??, November 14, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        185/.

Schmidt:2021:CSF

[Sch21] James R. Schmidt. CSVDataMerge: a simple and free program for concatenating experimental data files. Journal of Open Research Software, 9(1): 34–??, December 10, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        368/.

Siqueira:2016:PPP

[SdSS16] Abel Soares Siqueira, Raniere Costa da Silva, and Luiz-Rafael Santos. Perprof-py: A Python package for performance profile of mathematical optimization software. Journal of Open Research Software, 4(1):e12–??, April 22, 2016. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        81/.

Shakir:2017:PGG

[SGPHD^{+}17] Dzhoshkun Ismail Shakir, Luis Carlos García-Peraza-Herrera, Pankaj Daga, Tom Doel, Matthew J. Clarkson, Sébastien Ourselin, and Tom Vercauteren. GIFT-Grab: Real-time C++ and Python multi-channel video capture, processing and encoding API. Journal of Open Research Software, 5(1):27–??, October 09, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
169/.

Silva:2014:OST

[SM14a] Ikaro Silva and George Moody.

An open-source toolbox for analysing and processing PhysioNet databases in MATLAB and Octave. Journal of Open Research Software, 2(1):e27– ??, September 24, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bi/.

Stodden:2014:BPC

[SM14b] Victoria Stodden and Sheila Miguez. Best practices for computational science: Software infrastructure and environments for reproducible and extensible research. Journal of Open Research Software, 2(1):e21–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        ay/.

Sahlmann:2020:GGU

[SM20] Lisa Sahlmann and Christian B. Mendl. GuiTeNet: a graphical user interface for tensor networks. Journal of Open Research Software, 8(1):29–??, December 15, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        304/.

Sochat:2022:RSE

[SMC^{+}22] Vanessa Sochat, Nicholas May, Ian Cosden, Carlos Martinez-Ortiz, and Sadie Bartholomew. The Research Software Encyclopedia: a community framework to define research software. Journal of Open Research Software, 10(1):??, March 04, 2022. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        359/.

Stieg:2019:CSS

[SMG19] Julian Stieg, Peter Marks, and Lasse Gerrits. UN-CODE: Software for structuring and visualizing collective decisionmaking based on qualitative data. Journal of Open Research Software, 7 (1):25–??, July 31, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
246/.

Sochat:2023:ADC

[SMS^{+}23] Vanessa Sochat, Matthieu Muffato, Audrey Stott, Marco De La Pierre, and Georgia Stuart. Automated discovery of container executables. Journal of Open Research Software, 11(1):??, ???? 2023. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
451.

Sulzer:2021:PBM

[SMT^{+}21] Valentin Sulzer, Scott G. Marquis, Robert Timms, Martin Robinson, and S. Jon Chapman. Python battery mathematical modelling (PyBaMM). Journal of Open Research Software, 9(1):14–??, June 08, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
309/.

Slaughter:2015:CIC

[SPG^{+}15] Andrew E. Slaughter, John W.

Peterson, Derek R. Gaston, Cody J. Permann, David Andrs, and Jason M. Miller. Continuous integration for concurrent MOOSE framework and application development on GitHub. Journal of Open Research Software, 3(1):e14–??, November 20, 2015. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        bx/.

Staal:2020:GMM

[SR20] Tobias Stål and Anya M. Reading. A grid for multidimensional and multivariate spatial representation and data processing. Journal of Open Research Software, 8(1):2–??, January 30, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        287/.

Saavedra:2018:OMI

[SS18] Pablo Saavedra and Clemens Simmer. An Octave/MATLAB(R) interface for rapid processing of SMOS L1C full polarization brightness temperature. Journal of Open Research Software, 6(1):2–??, January 08, 2018. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        165/.

Slugocki:2019:BTM

[SSB19] Michael Slugocki, Allison B. Sekuler, and Patrick Bennett. BayesFit: a tool for modeling psychophysical data using Bayesian inference. Journal of Open Research Software, 7(1):2–??, January 17, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
202/.

Stone:2015:WMT

[ST15] James Stone and John Towse.

A working memory test battery: Java-based collection of seven working memory tasks. Journal of Open Research Software, 3(1):e5–??, June 05, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
br/.

Stadler:2021:PPB

[Sta21] Konstantin Stadler. Pymrio — a Python based multi-regional input-output analysis toolbox. Journal of Open Research Software, 9(1):8–??, May 11, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
251/.

Stock:2016:OSS

[Sto16] Andy Stock. Open source software for mapping human impacts on marine ecosystems with an additive model. Journal of Open Research Software, 4(1):e21–??, June 07, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Seibold:2019:MRP

[SZH19] Heidi Seibold, Achim Zeileis, and Torsten Hothorn. model4you: an R package for personalised treatment effect estimation. Journal of Open Research Software, 7(1):17–??, May 15, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        219/.

Suriarachchi:2015:PKC

[SZP15] Isuru Suriarachchi, Quan Zhou, and Beth Plale. Komadu: A capture and visualization system for scientific data provenance. Journal of Open Research Software, 3(1):e4–??, March 30, 2015. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        bq/.

Tauscher:2018:PBP

[Tau18]  Helga Tauscher.  Billie.

a prototypical framework for building information model visualization. Journal of Open Research Software, 6 (1):18–??, May 18, 2018. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        206/.

Tong:2022:GAI

[TCD^{+}22] Xin Tong, Sou-Cheng T. Choi, Yuhan Ding, Fred J. Hickernell, Lan Jiang, Lluís Antoni Jiménez Rugama, Jagadeeswaran Rathinavel, Kan Zhang, Yizhi Zhang, and Xuan

Zhou. Guaranteed Automatic
Integration Library (GAIL):

an open-source MATLAB library for function approximation, optimization, and integration. Journal of Open Research Software, 10(1):??, July 29, 2022. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
381/.

Trainer:2014:BES

[TCH14] Erik Trainer, Chalalai Chaihirunkarn, and James Herbsleb. The big effects of short-term efforts: Mentorship and code integration in open source scientific software. Journal of Open Research Software, 2(1):e18–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bc/.

Thompson:2020:SUA

[TDX^{+}20] Stephen Thompson, Thomas Dowrick, Goufang Xiao, João Ramalhinho, Maria Robu, Mian Ahmad, Dan Taylor, and Matthew J. Clarkson. SnappySonic: an ultrasound acquisition replay simulator. Journal of Open Research Software, 8(1):8–??, March 30, 2020. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        289/.

Turner:2021:OLE

[TLR21] Ross J. Turner, Rebecca B. Latto, and Anya M. Reading. An ObsPy library for event detection and seismic attribute calculation: Preparing waveforms for automated analysis. Journal of Open Research Software, 9(1):29–??, October 19, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        365/.

Thielicke:2014:PPT

[TS14] William Thielicke and Eize Stamhuis. PIVlab — towards user-friendly, affordable and accurate digital particle image velocimetry in MATLAB. Journal of Open Research Software, 2(1):e30–??, October 16, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        bl/.

Thielicke:2021:PIV

[TS21] William Thielicke and René Sonntag. Particle image velocimetry for MATLAB: Accuracy and enhanced algorithms in PIVlab. Journal of Open Research Software, 9(1):12–??, May 31, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.

com/articles/10.5334/jors.
334/.

Usher:2019:SFI

[UR19] Will Usher and Tom Russell. A software framework for the integration of infrastructure simulation models. Journal of Open Research Software, 7(1):16–??, May 07, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
265/.

Vousden:2016:VMF

[VBAF16] Mark Vousden, Marc-Antonio Bisotti, Maximilian Albert, and Hans Fangohr. Virtual micromagnetics: A framework for accessible and reproducible micromagnetic simulation. Journal of Open Research Software, 4(1):e41–??, October 31, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
141/.

vandenBerg:2019:OTP

[vdB19] Ewout van den Berg. The Ocean Tensor package. Journal of Open Research Software, 7(1):26–??, August 01, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Varner:2019:WPP

[VEV^{+}19] James F. Varner, Noor Eldabagh, Derek Volta, Reem Eldabagh, and Jonathan J. Foley IV. WPTherml: a Python package for the design of materials for harnessing heat. Journal of Open Research Software, 7(1):28–??, August 19, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
271/.

vanGent:2019:AND

[vGFvNvA19] Paul van Gent, Haneen Farah, Nicole van Nes, and Bart van Arem. Analysing noisy driver physiology real-time using offthe-shelf sensors: Heart rate analysis software from the Taking the Fast Lane Project. Journal of Open Research Software, 7(1):32–??, October 29, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
241/.

vanGompel:2016:EGS

[vGvdB16] Maarten van Gompel and Antal van den Bosch. Efficient ngram, skipgram and flexgram modelling with Colibri Core. Journal of Open Research Software, 4(1):e30–??, August 02, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
105/.

Vernon:2019:GHF

[VHT^{+}19] Chris R. Vernon, Mohamad I.

Hejazi, Sean W. D. Turner, Yaling Liu, Caleb J. Braun, Xinya Li, and Robert P. Link. A global hydrologic framework to accelerate scientific discovery. Journal of Open Research Software, 7(1):1–??, January 07, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
245/.

Vernon:2018:PDL

[VLC^{+}18] Chris R. Vernon, Yannick Le Page, Min Chen, Maoyi Huang, Katherine V. Calvin, Ian P. Kraucunas, and Caleb J. Braun. Demeter — a land use and land cover change disaggregation model. Journal of Open Research Software, 6(1):15–??, April 20, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
208/.

Venters:2014:BME

[VLG^{+}14] Colin Venters, Lydia Lau, Michael Griffiths, Violeta Holmes, Rupert Ward, Caroline Jay, Charlie Dibsdale, and Jie Xu. The blind men and the elephant: Towards an empirical evaluation framework for software sustainability. Journal of Open Research Software, 2(1):e8–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        ao/.

VanWaes:2019:MCT

[VLPV19] Luuk Van Waes, Mariëlle Leijten, Tom Pauwaert, and Eric Van Horenbeeck. A multilingual copy task: Measuring typing and motor skills in writing with Inputlog. Journal of Open Research Software, 7(1):30–??, October 16, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        234/.

Vernon:2018:PCG

[VZR^{+}18] Chris R. Vernon, Nino Zuljevic, Jennie S. Rice, Timothy E. Seiple, Michael C. W. Kintner-Meyer, Nathalie Voisin, Ian P. Kraucunas, Jin Chunlian, Jarrod Olson, Laurel Schmidt, Scott L. Morris, and Pralit Patel. CERF — a geospatial model for assessing future energy production technology expansion feasibility. Journal of Open Research Software, 6(1):20–??, August 06, 2018. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        227/.

Wagenaar:2017:PVD

[Wag17] Daniel A. Wagenaar. VScope — data acquisition and analysis for voltage-sensitive dye imaging using multiple cameras and electrophysiology. Journal of Open Research Software, 5(1):23–??, September 13, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
176/.

Wijnen:2016:FOS

[WAH^{+}16] Bas Wijnen, Gerald C. Anzalone, Amberlee S. Haselhuhn, P. G. Sanders, and Joshua M. Pearce. Free and open-source control software for 3-D motion and processing. Journal of Open Research Software, 4(1):e2–??, January 27, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
78/.

Wang:2019:OSS

[Wan19]  Y. Q. Wang.  An open

source software suite for multi-dimensional meteorological data computation and visualisation. Journal of Open Research Software, 7 (1):21–??, July 11, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
267/.

Wethor:2022:STS

[WH22] Gabi Wethor and Matthew L.

Hale.  StudySandboxx: a

tool for scraping, sandboxing, preserving, and preparing interactive Web sites for use in human-computer interaction and behavioral studies. Journal of Open Research Software, 10(1):??, July 12, 2022. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
274/.

Walker:2019:ATS

[WKC19] Paul Walker, Ulrich Krohn, and David Carty. ARBTools: a tricubic spline interpolator for three-dimensional scalar or vector fields. Journal of Open Research Software, 7(1):12–??, April 18, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
258/.

Wheeler:2019:PPF

[WKD^{+}19] Daniel Wheeler, Trevor Keller, Stephen J. DeWitt, Andrea M. Jokisaari, Daniel Schwen, Jonathan E. Guyer, Larry K. Aagesen, Olle G. Heinonen, Michael R. Tonks, Peter W. Voorhees, and James A. Warren. PFHub: The phasefield community hub. Journal of Open Research Software, 7(1):29–??, September 24, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
276/.

Wagner:2013:PII

[WL13] Thorsten Wagner and Hans-Gerd Lipinski. IJBlob: An ImageJ library for connected component analysis and shape analysis. Journal of Open Research Software, 1(1):e6–??, October 14, 2013. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
ae/.

Williams:2020:PSP

[WL20] Brendan Williams and Michael Lindner. pyfMRIqc: a software package for raw fMRI data quality assurance. Journal of Open Research Software, 8(1):23–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
280/.

Wild:2019:SRB

[WLA19] Thomas B. Wild, Daniel P.

Loucks, and George W. Annandale. SedSim: a river basin simulation screening model for reservoir management of sediment, water, and hydropower. Journal of Open Research Software, 7(1):22–??, July 17, 2019. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.

Wright:2015:PDP

[WS15] John G. Wright and B. Sriram Shastry. DiracQ: A package for algebraic manipulation of noncommuting quantum variables. Journal of Open Research Software, 3(1):e13–??, November 20, 2015. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        cb/.

White:2019:DJR

[WTLB19] Lyndon White, Roberto Togneri, Wei Liu, and Mohammed Ben-

                                        namoun. DataDeps.jl: Re-

peatable data setup for reproducible data science. Journal of Open Research Software, 7(1):33–??, October 29, 2019. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        244/.

Windecker:2021:MRP

[WVTTG21] Saras M. Windecker, Peter A. Vesk, Stacey M. Trevathan-Tackett, and Nick Golding. mixchar: an R package for the deconvolution of thermal decay curves. Journal of Open Research Software, 9(1):27–??, October 01, 2021. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        249/.

Xiong:2016:PKT

[XXX16] Hang Xiong, Pin Xiong, and Hui Xiong. KAMG: A tool for converting blood ties and affinity ties into adjacency matrices. Journal of Open Research Software, 4(1):e28–??, July 19, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
99/.

Yu:2017:PMR

[YL17] Qingzhao Yu and Bin Li. mma:

An R package for mediation analysis with multiple mediators. Journal of Open Research Software, 5(1):11–??, April 12, 2017. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
160/.

Zeitvogel:2016:PSI

[ZO16] Fabian Zeitvogel and Martin Obst. ScatterJn: An ImageJ plugin for scatterplotmatrix analysis and classification of spatially resolved analytical microscopy data. Journal of Open Research Software, 4(1):e5–??, February 03, 2016. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
89/.

Zentner:2014:NOE

[ZZF^{+}14] Lynn Zentner, Michael Zentner, Victoria Farnsworth, Michael McLennan, Krishna Madhavan, and Gerhard Klimeck. nanoHUB.org: Experiences and challenges in software sustainability for a large scientific community. Journal of Open Research Software, 2(1):e19–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        bd/.